### v2

# Argumentative Experience: Reducing Confirmation Bias on Controversial Issues through LLM-Generated Multi-Persona Debates

Abstract.
Large language models (LLMs) are enabling designers to give life to exciting new user experiences for information access. In this work, we present a system that generates LLM personas to debate a topic of interest from different perspectives. How might information seekers use and benefit from such a system? Can centering information access around diverse viewpoints help to mitigate thorny challenges like confirmation bias in which information seekers over-trust search results matching existing beliefs? How do potential biases and hallucinations in LLMs play out alongside human users who are also fallible and possibly biased?
Our study exposes participants to multiple viewpoints on controversial issues via a mixed-methods, within-subjects study. We use eye-tracking metrics to quantitatively assess cognitive engagement alongside qualitative feedback. Compared to a baseline search system, we see more creative interactions and diverse information-seeking with our multi-persona debate system, which more effectively reduces user confirmation bias and conviction toward their initial beliefs. Overall, our study contributes to the emerging design space of LLM-based information access systems, specifically investigating the potential of simulated personas to promote greater exposure to information diversity, emulate collective intelligence, and mitigate bias in information seeking.
Large Language Models (LLMs), Confirmation bias, Multi-persona debate system, User-centered design, Eye-tracking experiment
copyright: noneccs: Human-centered computing User studiesccs: Human-centered computing User centered design
1. Introduction
Large language models (LLMs) offer a novel design space to explore new ways to support human information seeking. For instance, commercial search engines and services like ChatGPT and Perplexity can synthesize information from multiple sources to provide direct answers to user questions [^1] ChatGPT: https://openai.com/index/chatgpt/, Perplexity: https://www.perplexity.ai/.. Recent work has also explored leveraging the internal knowledge and reasoning capabilities of LLMs to simulate human personas with diverse perspectives (Prpa et al., 2024; Park et al., 2022; Chen et al., 2024; Khan et al., 2024). Imagine a user could generate a set of LLM personas on-demand to debate any topic of interest from multiple perspectives. How should we design such a system to maximize potential benefit to information seekers (e.g., information diversity) while also minimizing associated risks (e.g., LLM hallucination and biases interacting with human biases)?
In traditional information-seeking process, users retrieve facts, evidence, and other information from various sources, often guided by their pre-existing beliefs. This can lead to confirmation bias (Berthet, 2021), where users favor information that supports their prior assumptions. Such biases are prevalent not only in standard information access systems, like search engines, but also in conversational interactions, where users actively shape the direction of the discourse (Sharma et al., 2024; Khan et al., 2024). However, LLM-powered diverse information-seeking could be intentionally designed to provide broad coverage and present contrasting views, regardless of prior user beliefs. Recent design explorations of multi-persona systems have been encouraging, with subjective or indirect user feedback suggesting mitigated confirmation bias (Park et al., 2023a; Chiang et al., 2024; Zhang et al., 2024).
In this work, we explore a novel information seeking experience: letting users launch a debate on any topic of interest in which generated LLM personas present diverse perspectives on the topic. We explore the active and autonomous nature of LLM agents in this setting as they present information and arguments, creating an argumentative experience between humans and AI. In focusing the user experience on exposure to diverse perspectives, we particularly investigated whether exposure to diverse information via generated personas could help reduce user confirmation bias.
Participants in our study launched LLM debates on controversial topics spanning social issues (e.g., ”Should abortion bans be illegal?”) to imaginary scenarios (e.g., ”Should space be colonized?”) We first assessed initial user attitudes toward these topics, then asked users to engage with a baseline search system and our multi-persona debate system. We used a mixed-methods approach to capture participants’ cognitive perceptions, such as real-time eye-tracking metrics, behavioral data, such as user interactions, and qualitative feedback through retrospective think-aloud sessions. These data not only help assess the relative performance of our multi-persona debate system, but also address limitations of other recent exploratory studies on LLM-persona systems, which primarily relied on self-reported metrics and subjective feedback (Park et al., 2023a; Zhang et al., 2024). Our study thus complements this prior work in moving toward a more comprehensive view of how use of LLM personas can impact user information consumption.
Our study investigates three key research questions:
RQ1:
What is the overall user experience when conceptualizing controversial topics through a multi-persona debate interaction simulated by LLM?
RQ2:
Compared to the retrieval-based system, how effective is the multi-persona debate system in mitigating confirmation bias?
RQ3:
Compared to the retrieval-based system, how does the multi-persona debate system finally influence user pre-existing beliefs?
Through a combined analysis of both quantitative and qualitative data, our findings reveal that the argumentative experience powered by LLM personas helps foster constructive deliberation on contrasting viewpoints, making individuals more likely to consume content that challenges their pre-existing beliefs. This happens because persona-based debates present information in a more dynamic, appealing, and persuasive manner, reducing user cognitive load and enhancing interactivity so that people are willing to engage with challenging content. In summary, our work investigates persona-based debate systems to gain deeper insights into human cognitive perceptions and behavioral patterns when interacting with them. We assess the effectiveness of these systems in reducing confirmation bias during information seeking and offer broader design implications to guide the future development of persona-based systems.
2. Related Work
2.1. The Social Simulation and Role-Play of LLM
Over the past five years, LLMs (e.g., GPT, Llama, Gemini, etc.) have rapidly advanced and today represent the state-of-the-art in performance across various natural language tasks [^2] GPT: https://openai.com/index/gpt-4/, Llama: https://www.llama.com/, Gemini: https://gemini.google.com/. Additionally, by being trained on the increasing amounts of web data and computational resources (Kaplan et al., 2020), LLMs have developed strong reasoning and association abilities, as well as internal knowledge that captures social knowledge and cultural norms.
To instruct and tailor content generated by LLMs, prompting is a specific technique used by practitioners, providing initial inputs or cues to a language model (Liu et al., 2023). Researchers have explored creative ways to assess and extend the human-like capabilities of LLMs through prompting (Wang et al., 2021; Wei et al., 2022; Yao et al., 2022). One novel application is the development of agent-based systems where LLMs are used to create highly autonomous agents. LLM personas (or Social Simulacra) are a prominent example. Specifically, by incorporating different beliefs, desires, goals, and characteristics into the prompts, LLMs can role-play to simulate human reactions and behaviors (Shanahan et al., 2023; Park et al., 2022; Chen et al., 2024; Tseng et al., 2024). This approach offers new opportunities for enhancing user experiences in Computer Supported Cooperative Work (CSCW) and other areas of Human-Computer Interaction (HCI). In these fields, generative agents can provide more dynamic, interactive, and autonomous user experiences (Park et al., 2023b). Researchers have explored different use cases, such as using social simulacra to conduct user studies (Argyle et al., 2022; Park et al., 2022), creating generative agents in an interactive sandbox environment (Park et al., 2023b), or directly using LLMs as judges to assist decision-making tasks (Pan et al., 2024), ranging from data annotation (Tan et al., 2024) to language evaluation (Wang et al., 2023).
However, researchers also highlight important challenges in the real-world experiments that employ LLM as the core technique to design persona systems, as well as agent-based systems in general. For example, in particular on the development of social simulacra, LLMs can reflect polarized human opinions regarding different social issues, including misinformation (Neumann et al., 2024), climate change (Santurkar et al., 2023), and politics (Taubenfeld et al., 2024). More importantly, the broad idea of agent-based systems also raises important ethical questions about the idea of replacing humans (Prpa et al., 2024). In summary, the development and implementation of LLM-based systems are still in their early stages. Ongoing research is needed to investigate both the advantages and challenges of applying these systems to different real-world scenarios.
2.2. Confirmation Biases in Information Seeking
Confirmation bias is one of the cognitive biases (Berthet, 2021) which refers to the human tendency to seek and interpret evidence that supports their existing beliefs, expectations, or hypothesis (Nickerson, 1998), regardless of the plausibility of those beliefs (Edwards and Smith, 1996). Confirmation bias consists of two aspects: the cognitive aspect (disposition) and the motivational aspect (manifestation) (Berthet et al., 2022). The cognitive aspect of confirmation bias focuses on human tendencies (Berthet et al., 2022; Baron, 2023). It represents a stable psychological disposition that reflects an individual’s tendency toward confirmatory information seeking and evaluation, which can be reliably measured through validated psychometric instruments (Ardi and Pradiri, 2021; Rassin, 2008; Berthet, 2021). The motivational aspect focuses on the relationship with users’ prior beliefs (Baron, 1995). This aspect relates to users’ pre-existing beliefs within specific contexts and manifests dynamically during information-seeking activities, where individuals tend to select information aligning with their attitudes, allocate more attention to confirming content, and assign greater weight to belief-consistent evidence during evaluation (Vedejová and Čavojová, 2022).
Previous research has shown that confirmation bias is widely manifested in seeking behavior, influencing how users interact with information. Specifically, users may be more likely to select and spend more time reading attitude-consistent results (White, 2013; Suzuki and Yamamoto, 2020), while avoiding information that challenges their prior beliefs (El and Moussaoui, 2022; Knobloch-Westerwick et al., 2015; Dickinson, 2020). This behavior may further result in filter bubble and echo chamber (Quattrociocchi, 2017), where like-minded people gather, communicate and interact as an identified community, reinforcing shared beliefs and perspectives (Garrett, 2017).
To mitigate such biases, HCI and CSCW scholars have explored design strategies in visualization and recommendation systems to encourage user engagement with diverse content. One established approach is to increase exposure to opposing viewpoints. For instance, Lex et al. (2018) mixed tweets and hashtags from various partisan groups, Schweiger et al. (2014) used tag clouds to display opposing viewpoints, and Huang et al. (2012) presented counterarguments on a separate screen. Another approach is to increase the system interactivity and user engagement. For instance, Sude et al. (2021) added a ”voting” feature for search results, and Rieger et al. (2021) applied warning labels like ”this search result may reinforce your opinion,” requiring users to click a button to view them. These features help users to be more aware of their information seeking and decision framing process. Additionally, developers created search systems to balance different search results by learning user existing information activities, such as Schwind et al. (2012)’s search system that highlights search result inconsistent with user preferences reflected on their search queries and Jasim et al. (2022)’s bias mitigation recommendation model that suggests diverse sentiment review for online products, informed by user past browsing history. Aforementioned studies suggest that presenting diverse information in a more vivid way can motivate users to engage with content that differs from their prior views.
However, it’s important to recognize that these design and system enhancements remain centered on a static, passive model of information consumption, characteristic of traditional search contexts where information is retrieved and displayed rather than dynamically presented. While increasing exposure to diverse information through external cognitive forcing functions can increase information diversity, this approach may not inherently enhance user engagement or persuasiveness. In contrast, LLM-powered information seeking presents a novel interaction paradigm that enables more diverse, interactive, and dynamic engagement, potentially reducing confirmation bias.
2.3. Opportunities of LLM Constructive Deliberation on Contrasting Viewpoints
Confirmation bias is often associated with polarization effects in which social divides serve to establish and/or reinforce certain viewpoints. While exposure to opposing views in a hostile setting can heighten polarization (Bail et al., 2018), constructive deliberation between opposing views can actually promote empathy, tolerance, and sharing ideas, and reduce polarization (Mutz, 2002; Porter and Schumann, 2018; Levendusky and Stecula, 2021; Bächtiger and Parkinson, 2019; Kim et al., 2021).
Researchers have explored how LLMs can facilitate constructive deliberation and enhance engagement with contrasting viewpoints through mediated discourse.
From the technical side, Tessler et al. (2024) develop the Habermas Machine, a system of LLMs based on Chinchilla which generates statements that express common ground amongst a set of opposing views. Kambhatla et al. (2024) operationalize receptiveness into a computational framework, and use an LLM in combination with the framework to reframe Reddit comments to be more receptive to other users’ comments. Other work also studies grounding, or constructing a shared basis of understanding, in promoting positive discussions on opposing views by fine-tuning specific language model (Cho and May, 2020; Shaikh et al., 2024b).
Researchers in social science and HCI also design LLM-powered tools to facilitate constructive deliberation. For example, Argyle et al. (2023) aim to improve conversation quality in divisive online political conversations using an AI chat assistant. Similarly, Zhang et al. (2024) designed a multi-agent dialogue system in social media conversations. Although no comparison was made with other baseline conditions, they found that participants were influenced either by their own interests or by the system itself to actively explore diverse perspectives, helping to reduce their filter bubbles. Chiang et al. (2024) directly use LLM to play devils’ advocate in AI-assisted group decision-making and found that participants were more likely to form an appropriate trust to adopt AI decisions.
Overall, we see LLMs increasingly being used to reduce polarization and intolerance in ways that might complement LLM-tooling to mitigate related biases. Notably, LLM could be designed to simulate personas that take more critical stances according to predefined roles and characters, which might challenge user prior beliefs. By presenting a new user experience built around diverse feedback and engagements than traditional search, this approach might help counteract the biased information-seeking behaviors commonly seen in traditional search systems.
Research Focus of this Study. Building on prior research, we aim to thoroughly and objectively assess the effectiveness of LLM-generated personas in reducing confirmation bias within an automated debate context. LLM-powered diverse information-seeking via personas demonstrates how LLM agents actively and autonomously can present information and arguments, creating an argumentative experience between humans and AI. Unlike traditional information-seeking that is predominantly user-driven, this argumentative experience could foster a more interactive and collaborative exchange of information to reduce both motivational and cognitive confirmation bias.
While prior studies have examined LLM-generated personas for product recommendations (Park et al., 2023a), reducing filter bubbles through qualitative feedback (Zhang et al., 2024), and generating counterarguments in group decision-making (Chiang et al., 2024), our study objective is to ground the hypotheses (as highlighted above) from a psychological cognitive perspective.
We designed a multi-persona debate system that presents diverse viewpoints through argumentative discourse on controversial issues. By comparing this system with a baseline retrieval-based approach that presents balanced content snippets from opposing perspectives, we evaluate how dynamic, argumentative information presentation influences confirmation bias and belief formation. Through eye-tracking analysis capturing users’ involuntary (bottom-up) and goal-directed (top-down) eye movements, we examine how this active, LLM-driven interaction paradigm shapes user engagement.
In the following sections, we present the system design for the multi-persona debate system (Section 3) and detail the study procedure (Section 4).
3. System Design
This section describes the design and development of our multi-persona debate system, which serves as the experimental interface for our study. As an overview of the multi-persona debate system and how users may interact with it, we present a simplified user-system workflow in Figure 1. Details are presented regarding the envisioned usage scenario (Section 3.1), design objectives (Section 3.2), interface features (Section 3.3), and system architecture (Section 3.4).
Figure 1. Overview of the Multi-Persona Debate System
3.1. Envisioned Use Case
Controversial social issues, such as gun control, abortion, climate change, etc., often attract heightened media attention and public interest, where partisanship and political divisions can influence adopted stances. This suggests value in exploring new ways by which people might navigate complex and nuanced issues of divergent positions and ideologies. We assume users are at least somewhat curious and interested to explore diverse views (or they would not be using our system). Someone particularly firm their own beliefs may be curious about how anyone else could think otherwise.
Our multi-persona debate system offers a novel form of information access for such topics, using LLM personas to generate diverse and, possibly, contrasting perspectives. A user begins by entering a topic of interest, similar in spirit to a traditional search query. The system then generates suggested personas to provide broad coverage of differing views. The user can freely keep these generated personas, revise them, or add others. Following this, the user initiates as many rounds of debate as they find valuable, further revising the set of personas at any time to explore additional perspectives. By simulating a debate between personas, each advocating for a distinct perspective, users are exposed to a rich array of arguments, counterarguments, and the complex interplay of ideas that characterize real-world discussions.
For example, when debating a topic such as ”Should space be colonized?”, the system presents perspectives from diverse stakeholders such as scientists, environmentalists, and economists, each offering unique insights on the matter. In the debate, personas generated by LLMs present their arguments in a conversational setting. A simulated “space scientist” may argue that colonizing space is crucial for humanity’s survival, while a simulated “environmentalist” might oppose this stance, raising concerns about the potential environmental and ethical implications.
As users observe the debate, some arguments may naturally resonate with them more than others. To promote more active engagement and information retention, users can bookmark any argument of particular interest they would like to review later, as well as to record their agreement or disagreement with it (via thumbs up or down buttons) to remind themselves later when they review bookmarked content. Users can review their bookmarked content at any time during or after the debate. This interactive experience – customizing the debate and bookmarking arguments of interest – seeks to help users organize and track the main points of each perspective in order to foster a deeper, more critical engagement with the topic at hand.
3.2. Design Objectives
Motivated by the aforementioned usage scenario, our design goals are as follows:
  1. (1)
    Broad Representation: We generate LLM personas that provide comprehensive coverage of major viewpoints on each topic. This development requires a deliberate and balanced approach to perspective generation to ensure balanced representation while maintaining a calibrated spectrum from opinion-based to factual discourse.
  2. (2)
    Dynamic Debate Interactions: We strive to create dynamic debate interactions among personas. This includes the ability for personas to respond to each other’s arguments, pose questions, and provide rebuttals, mirroring the fluid nature of human dialogue.
  3. (3)
    User Engagement and Interactivity: We seek to promote user engagement by enabling users to save salient responses, modify persona parameters, and initiate successive debate rounds. This interaction design fosters active participation in dialectical exchanges rather than passive observation.
Figure 2. The interface of our multi-persona debate system is shown for the topic “Should human colonize space?” LLM-generated personas are shown at left, which users can freely modify, replace, or delete. Once the debate starts, thumbs up/down buttons to the top-right of each argument can be used to bookmark it for the user’s later review, recording whether or not the given argument resonated with the user’s own thinking. Bookmarked arguments are collected for the user’s review in the Summary tab.
3.3. Interface Features
Figure 2 presents the user interface of our LLM debate system. We describe the major interactive features below.
3.3.1. Controversial Topics Entry or Selection
At the start, users can either enter a topic, possibly controversial, in the search bar or choose one from the list of suggested topics A. The system then begins generating LLM-personas relevant to that topic.
3.3.2. Persona Creation
As displayed in B, each persona is represented by a title and description, along with automatically assigned colors and icons to help users easily identify and differentiate between them (which emerged as a need during early pilot testing). We set a predefined number of personas at the beginning (N=5), but users are free to generate or manually create as many personas as they want. The specific prompt instructions (described in Section 3.4.1) aim to generate diverse LLM personas that represent a range of viewpoints and social roles on the selected topic. This approach aligns with our primary design goal of comprehensive representation.
3.3.3. Interactive Debate Flow
During the design of this interface, we addressed a critical design consideration regarding the optimal level of inter-persona interaction within each debate round.
In a structured human debate [^3] https://debatingmatters.com/debate-timing-structure/ where debaters are separated into two teams holding opposing views, responses may either focus on arguments presented by a specific debater or address all previously discussed points made by members from the opposing team. Additionally, arguments may unfold sequentially under the guidance of a moderator or occur simultaneously, with debaters from different teams engaging back-and-forth in animated exchanges.
Given our second design goal of fostering dynamic debate interaction while preventing the user from being overload, we have the personas engage in a round-robin (sequential) discussion. The debated content is displayed in a traditional conversational format C, making it easier for users to follow the debate. The prompt instructions (Section 3.4.2) encourage each persona to sequentially formulate their arguments by referring to the entire debate history.
A debate round concludes once all personas have (in order) presented their arguments. Afterward, users can select, deselect, add, or edit personas, and start a new round of debate by clicking the “Start Debate” button. Each time an additional round of debate is initiated, a new tab is created in the interface. This enables users to navigate back-and-forth across different rounds of the debate in reviewing the arguments presented. It is motivated by our third design goal of fostering user engagement and interactivity.
3.3.4. Bookmarks
With multiple personas presenting arguments over several rounds of debate, the number of arguments can quickly grow, which risks overwhelming the user. To help keep the set of arguments presented more manageable, we introduce bookmarking in our interface D, allowing the user to save arguments of particular note for later review. By clicking up/down thumb icons, users can further record whether or not a given argument resonated with their own thinking. These user bookmarks are collected and shown in the main interface’s ‘Summary’ tab, to the right of the latest debate round, as displayed in C. Users can review their bookmarked content at any time during or after the debate.
3.4. System Architecture
This section describes the architecture of our system, detailing the prompt instructions to support each interactive feature, including persona creation (Section 3.4.1) and debate generation (Section 3.4.2), language model choices (Section 3.4.3), and the web develop framework (Section 3.4.4).
3.4.1. Persona Creation
In creating personas for our multi-persona debate system, prior studies have pointed out several limitations inherent to LLMs (Park et al., 2022, 2023b), such as inconsistent persona generation and formatting. While LLMs can generate personas that clearly express their stances on controversial topics, determining how to effectively prompt the LLM to achieve this may not always be straightforward.
One problem is that assigning specific names to personas, real or fictional, can be problematic because readers may mistakenly believe the persona represents an actual person, as evidenced by prior studies that LLMs would inadvertently reveal real names from their training data (Huang et al., 2022). To mitigate misrepresentation and potential ethical concerns, we prompt the LLM to ”generate a title for the persona, not a name”, while enabling users to override this default setting through customization. In generating persona descriptions, the LLM may include detailed background information without explicitly clarifying the persona’s stance on the topic, yet explicitly linking persona background to specific stance risks perpetuating stereotypes or creating caricatures of those backgrounds. To address this, we first prompt the LLM to “describe background (instead of their stance)” and then separately prompt it to “start the debate by briefly presenting an argument in support of its stance on the topic.”
We generate an initial set of personas to cover key perspectives on the issue. To enhance perspectival diversity, we incorporate existing personas as context when generating additional ones. Our prompt design includes specific instructions to maintain consistency across persona professions and descriptions (see Appendix A for prompt details).
3.4.2. Debate Generation
To facilitate the debate, we developed two prompts: one to initiate the debate and another to continue it based on prior debate conversations (see Appendix A for prompts details). However, it is important to note that LLMs are constrained by a maximum input length, known as the ‘context window’ (Wang et al., 2024) after which they may start to lose track of earlier portions of the input. This limitation can be problematic in our system, as personas need to reference the entire debate history to provide meaningful responses.
To overcome this challenge, inspired by Park et al. (2023b)’s work, when the debate history exceeds the LLM’s context window, we employ an extractive summarizer that compresses the transcript to fit within the required length while preserving the key points of the discussion. Specifically, we use the BERTExtractiveSummarizer (Miller, 2019), which provides a good balance between summarization performance and latency. To make the responses more concise, readable, and suitable, we also include prompt instructions “Generate at most 200 words” and “Use bullet points (in markdown format)”. This helps to reduce verbose, unstructured LLM generations and mitigate information overload.
3.4.3. Large Language Model Used
While the landscape of LLMs is vast and rapidly changing, OpenAI’s GPT models have thus far remained at the forefront, with GPT-$4$GPT-4: https://openai.com/index/gpt-4/ being among the most capable LLMs today. Its 128K token context window also allows it to perform well in long range tasks like ours. We thus adopted GPT-4 for the study presented here, though we note our system architecture is largely agnostic and could easily support other LLMs as well.
3.4.4. Web Framework
We developed the system UI using Streamlit [^5] https://streamlit.io/, a popular machine learning web development framework that allows development with an emphasis on interactivity and ease of use. It also promotes rapid prototyping, which made it easy to iterate on the interface design based on user feedback or emerging requirements.
4. Methods
To explore how users might engage with the multi-persona debate system and the system effectiveness of reducing confirmation bias, we conducted a controlled, within-subjects study by comparing the multi-persona debate system with a baseline retrieval-based system. In this section, we describe our formal evaluation protocol, including the experimental procedure and tasks (Section 4.1), participant recruitment (Section 4.2), the selection of controversial topics (Section 4.3), and data collection and analysis.
4.1. Experimental Procedure and Tasks
Figure 3. Flowchart of the experimental procedure.
As illustrated in Figure  3, the within-subject experiment started with eye calibration and task instructions. Participants were then randomly assigned to one of two study blocks, Block A or Block B. In Block A, participants utilized the baseline system to engage with the topic. We adopted ArgumentSearch [^6] https://argumentsearch.com/ (Daxenberger et al., 2020) as the baseline for our A/B experiments. ArgumentSearch is a retrieval-based search engine that identifies articles from the open web and organizes content snippets according to their stance on the user query (detailed description is provided in Appendix C).
In Block B, participants interacted with our multi-persona debate system as the experimental condition to explore and generate debate content with the given topic. Each block included a training task to first familiarize participants with the assigned system. Within each block, participants completed four randomized trials assigned with different controversial topics. Once they completed the trials in one block, they moved on to the other block and used the alternate system, ensuring they interacted both systems.
Each trial consisted of three parts: (i) Pre-task Questions (ii) Exploring the given topic with the assigned interface, and (iii) Post-task Questions. In the pre-task questions, we asked participants to explicitly rate their topic familiarity, pre-existing belief (i.e., stance) to the topic, and the confidence for their stance. For each trial, participants were presented with a controversial issue to explore the interface. After participants completing the task, we again evaluate their current stance to the topic and the confidence score, similar to the pre-task questions.
Upon completion of all eight trials across both systems, participants were required to complete a post-test questionnaire to assess their propensity for confirmation bias. Following this, a semi-structured interview was conducted to gain insight into the users’ interactions with the systems, their perception, as well as the usability and feedback. Details of the pre-, post-task questionnaires and interview questions are presented in Appendix B. The entire experimental study took approximately 1.5 to 2 hours to complete. This included 60 to 80 minutes for the experimental portion and an additional 15 to 25 minutes for the interview.
4.2. Participant recruitment
The study was conducted in a usability lab at a university, with $N=40$ participants (19 females, 18 males, and 3 non-binary identity). Particpants were recruited using convenience sampling, and were aged between 18 and 34 years (view demographic details in Appendix D). They came from a variety of academic disciplines, ranging from applied sciences, natural sciences, engineering, to social sciences, bringing diverse educational backgrounds. Participants were pre-screened for native-level English familiarity, 20/20 vision (uncorrected or corrected). Upon completion of the study, each participant was compensated.
In the data analysis, a total of 156 trials, derived from 39 participants each completing 4 tasks within the multi-persona debate system, were utilized for behavioral (Section 5.1) and eye-tracking analysis (Section 5.2). One participant data point was excluded due to the missing eye-tracking data during collection. All 40 participants were included in the user belief change analysis (Section  5.3).
4.3. Selection of Controversial Topics
These topics were sourced from procon.org [^7] ProCon.org_2024 covered a diverse range of categories, including health, science and technology, economics, and entertainment. These topics are formulated as yes/no proposed actions, such as ”Should abortion bans be illegal?” or ”Should space be colonized?”. The topic itself thus embeds a “pro” stance supporting the action vs. the “con” stance against it. We select topic categories to cover a range of important issues with broad societal relevance. Each topic was selected to ensure active public discussion, involvement from various groups, reliable information from credible sources, and a balanced view of arguments, without needing much specialized knowledge (view topic details in Appendix 9). In each block, one topic was selected for each category, resulting in four tasks per block and a total of eight tasks.
4.4. Data Collection and Analysis
Data collection during the experimental procedure incorporated multiple sources: pre-task questions, user interaction logs, post-task questions, system evaluation questionnaires, and interview recordings. Eye-tracking data were gathered using the Tobii TX-300 eye-tracker and processed through Tobii Pro Lab  [^8] https://www.tobii.com/products/software/behavior-research-software/tobii-pro-lab, a commercial usability and eye-tracking software that enabled raw gaze data collection, user interaction recording, and preliminary data cleaning for subsequent analysis.”
Further details about our analysis are provided below.
4.4.1. RQ1: What is the overall user experience when conceptualizing controversial topics through a multi-persona debate interaction simulated by LLM?
To categorize user interaction patterns with our LLM debate system, we conducted a clustering analysis using the following behavioral data:
Debate rounds:
The average number of rounds of debate generated.
Persona selected (pro/con):
The average number of personas adopting pro/con stances over the course of the debate for a given topic.
Debate duration:
The average time (in seconds) duration for each round of debate.
New persona:
The average number of new personas generated for each round.
Persona edits:
The average number of personas edited by participants for each round.
Bookmark visits:
The average number of times participants visited the summary tab for each round.
Bookmark made:
The average number of thumbs (up or down) clicked for each round.
Following the initial clustering analysis, we performed thematic coding to summarize the qualitative feedback provided. Since we compare user experience between two different interfaces, our qualitative themes focus on factors influencing participant comparisons between the A/B systems used. We present these findings in Section 5.1.
4.4.2. RQ2: Compared to the retrieval-based system, how effective is the multi-persona debate system in mitigating confirmation bias?
A primary research objective is to evaluate the system’s effectiveness in mitigating confirmation bias in information behavior, particularly in reducing selective exposure to attitude-consistent information. Through eye-tracking data analysis, we examine engagement patterns with content that diverges from users’ prior beliefs. Drawing on confirmation bias literature (described in Section 2.2), we hypothesize that increased temporal engagement with reading opposing viewpoints may indicate reduced confirmation bias through enhanced exposure to differing perspectives.
Additionally, we incorporate two key factors in evaluating the above hypothesis based on our study design. First, we considered the impact of users’ general tendency toward confirmation bias as a psychological disposition (described in Section 2.2), to understand how this human factor may affect the system’s effectiveness. The second factor is user topic familiarity, since we intentionally selected different social topics (described in Section 4.3). We thus formulate two further sub-questions regarding how users’ eye movements might differ based on these factors.
RQ2.1:
How do varying levels of topic familiarity and confirmation bias tendency influence user engagement with content?
RQ2.2:
How does use of our LLM debate system vs. the baseline system impact user engagement with content? Does this effect differ based on user levels of topic familiarity and their confirmation bias tendency?
The user topic familiarity and confirmation bias tendency were measured in the pre-task questions. We also measured two types of eye-tracking metrics (ET) and defined areas of interest (AOIs) to understand user engagement with the content. These measures are defined as follows:
Topic familiarity:
Users are categorized into five levels of familiarity based on a 5-point Likert scale, as collected in the pre-survey.
Confirmation biases tendency:
Participant levels of confirmation bias were measured using Berthet (2021)’s questionnaire. The final collected scores ranged from 0.34 to 0.88, with a median of 0.6. To categorize participant confirmation bias tendencies, we divided the scores into three equal groups, classifying them into low, medium, and high.
Eye-tracking:
Two metrics, including total fixation duration and fixation counts, were utilized to understand user engagement. Fixations were filtered to include only good-quality fixations, lasting between 100ms and 2000ms. Additionally, both fixation duration and counts were normalized relative to the length of the content.
Areas of interests:
We also label AOIs to identify participant selective exposure to attitude-consistent versus attitude-inconsistent information. For example, if the content supports the given topic and the participant also holds a supportive view, then the content would be labeled as attitude-consistent information. Fixations falling into these AOIs represent how user engage with the attitude-consistent information. Table 1 presents how we annotate attitude-consistent and attitude-inconsistent AOIs based on user pre-task belief and the stance of the content.
Pre-task Belief of Topics
Very Supportive /
Somewhat Supportive
Very Opposed /
Somewhat Opposed
Content Stance pro / moderate pro consistent inconsistent
con / moderate con inconsistent consistent
Table 1. Belief consistency of areas of interest (AOIs) in relation to pre-task belief and content stance.
We present the results for RQ2 in Section 5.2, focusing on user engagement of the content.
4.4.3. RQ3: Compared to the retrieval-based system, how does the multi-persona debate system finally influence user pre-existing beliefs?
We adopt a similar approach as with RQ2. Correspondingly, we formulate two sub-research questions relevant to topic familiarity and user confirmation bias tendency:
RQ3.1:
How do different levels of topic familiarity and confirmation bias tendency influence user belief changes?
RQ3.2:
How does our LLM debate system, compared to the baseline system, impact user final belief changes? Additionally, does this effect differ based on user levels of topic familiarity and their confirmation bias tendency?
User pre- and post- belief were collected separately in the pre- and post-task questions. We calculated three types of belief changes, including non-directional, bi-directional, and weighted belief changes based on user confidence, defined as follows:
Belief change:
Users are categorized into two groups of belief changes — changed vs. unchanged — based on the difference of belief measurement for each topic. User belief is measured based on a 5-point Likert scale (i.e., very supportive, somehow supportive, neutral, somehow opposed, and very opposed), collected in the pre- and post-task questions.
Bi-directional belief change:
Users are categorized into three groups of belief changes — less supportive, more supportive, and unchanged — showing the directional belief change for each topic, as calculated based on the previous 5-point Likert scale.
Weighted belief change:
We calculate the stance shift weighted by the average confidence level before and after using each system (Equation 1) [^9] In prior behavioral studies by Norman and Price (2015), relative confidence of people’s subjective responses were evaluated using average measures. Building on this, we evaluate participants’ relative confidence in their beliefs about the topics by calculating the average confidence level before and after the intervention.. This score reflects both the direction and strength of the belief change. User belief confidence is also measured by a 5-point Likert scale in the pre- and post-task questions.
(1) $\frac{(\textnormal{Belief}_{\textnormal{after}}-\textnormal{Belief}_{\textnormal{before}})(\textnormal{Confidence}_{\textnormal{after}}+\textnormal{Confidence}_{\textnormal{before}})}{2}$
Results for RQ3 are detailed in Section 5.3, which discusses belief changes.
5. Results
5.1. RQ1: User Experience in Conceptualizing Controversial Topics through Multi-Persona Debate
From the behavioral data collected in the debate system (as described in Section  4.4), users generated an average of 1.865 debate rounds per topic. For each round, an average of 1.734 personas supporting the topic and 1.972 personas opposing the topic were selected. Users made minimal edits (0.187 personas), clicked thumbs up/down 2.628 times to bookmark content, and accessed the bookmark tab only 0.121 times on average. Full descriptive data are provided in Table  6. As we recognized notable variations in the behavior data among participants, we performed a clustering analysis to better understand users’ engagement patterns and their relationship to both the system design and the topics explored.
Measures User behavioral groups
Group 0 Group 1 Group 2
# Rounds of debate -0.125 0.381 -0.455
# Persona selected (pro) -0.096 0.320 -0.385
# Persona selected (cons) -0.108 -0.631 0.893
t Debate duration 0.062 -0.451 0.576
# New persona -0.155 -0.209 0.349
# Persona edits -0.047 -0.109 0.166
# Bookmark visits 2.091 -0.366 -0.443
# Bookmark made 0.788 -0.333 0.094
Table 2. K-means clustering for interactive behavior in multi-persona debate system. Each column represents a cluster identified by the K-means algorithm, while each column corresponds to users’ interactions used for clustering. Values reflect the Z-scores of user behavioral measures across three behavioral groups. The color intensity reflects the magnitude of user engagement with each feature, with dark orange indicating higher values and lighter shades representing medium to low engagement.
We first applied K-means clustering to categorize participant behaviors while they used the debate system. Recognizing that participants might display different usage patterns across various topics, we treated each participant’s interaction with each topic in the system as a separate unit of analysis. This resulted in a total of 156 units of analysis, since 39 participants each engaged with 4 topics. In the K-means clustering process, features were standardized using Z-scores to normalize the data, ensuring that each feature contributed equally to the clustering algorithm and improving the comparability across different scales. As shown in Table 2, the clustering result yielded three distinct groups with an $R^{2}$ value of 0.255 and silhouette coefficient of 0.170.
The smallest cluster (N = 25, Group 0) was characterized by their exceptionally high scores in bookmark tab visits (Z = 2.091) and thumb clicks (Z = 0.788), while showing average usage in other behavior measures. Group 1 (N = 75) exhibited positive scores in the number of debate rounds (Z = 0.381) and the selection of supportive personas (Z = 0.320), but notable negative scores in debate duration (Z = -0.451) and the selection of opposed personas (Z = -0.631). Group 2 (N = 56) was distinguished by high scores in debate duration (Z = 0.576) and the selection of opposed personas (Z = 0.893). This group also showed moderate positive scores in new personas generated (Z = 0.349) and personas edits (Z = 0.166). By carefully reviewing the participant screen recordings and qualitative feedback during their use of the debate system and correlating these understandings with the clustering statistics, we identified and conceptualized three distinct types of usage behaviors. We characterize a usage pattern for each group as follows:
5.1.1. Active observers (Group 0)
Participants in this group frequently used thumbs up / down features to express their opinions while observing the debate, and often reviewed the bookmarked content. They typically engaged in more than two debate rounds and showed moderate involvement in selecting and editing personas, balancing supportive and opposing viewpoints. For example, during the debate topic of “Should humans colonize space?”, P14 remarked “I think there’s a lot of good points made. Mainly it comes down to the tangibility of it and the fact that it kind of does feel like we’d be abandoning the mess that we made here [the Earth].” Because P14 found it feasible for space colonization, this participant often gave thumbs-up to compelling arguments and thumbs-down to weaker ones. Similarly, others in this group proactively express their opinions using the thumbs features and spent more time on the bookmarked tabs. We see this pattern reflected in our envisioned use case (Section 3.1). It is marked by frequent use of feedback features, revisiting bookmarked content, and a balanced consideration of both supportive and opposing perspectives.
5.1.2. Multi-round explorers (Group 1)
Participants in this group were inclined to generate more rounds of debate (up to five) but spend less time reading and digesting arguments presented. Notably, we observed a preference for selecting personas that agreed with the given topic. From qualitative feedback, it appears that this may stem from the uncertainty about the topic. By initially choosing supportive personas, participants began to familiarize themselves with the subject matter. For example, in the debate over “Should humans colonize space?” P17 admitted a lack of knowledge in that area. The participant first selected two supportive personas, an aerospace engineer and a resource economist, to gain insights into the potential benefits of space colonization. As the debate progressed, P17 gradually included personas with different perspectives. This pattern reflects a strategy of building understanding through repetition and confirmation, with a gradual move toward considering alternative viewpoints as participant knowledge grows throughout the debate.
5.1.3. Critical readers (Group 2)
Participants in this group exhibited the highest level of engagement with the system. Compared to other two groups, they experienced the longest debate durations, selected more new personas, and frequently edited these personas. A notable characteristic was to actively seek out personas that disagreed with the topic. These behaviors suggest significant mental effort invested in generating, evaluating, refining personas, and reading content to gain a comprehensive understanding of the topic. This behavior might stem from participant propensity for critical thinking, as indicated from their qualitative interviews. For example, in the debate topic of “Should abortion be legal?” P14 initially supported the legalization of abortion, felt that they hadn’t explored the topic deeply enough “[Although] I’m kind of familiar, but I haven’t really delved in deep to all the different sides so I don’t want to take a side on it [quickly].” Despite holding a pro stance, this participant sought out opposite opinions to gain a more comprehensive view of the controversial issue. This pattern exemplifies how some participants, even with initial stances, actively engage with conflicting perspectives to ensure a balanced understanding of the topic.
In addition to identifying three general behavioral patterns, we found that topic-specific factors influence participants’ engagement with the debate system. Participants were also attentive to the expertise and backgrounds of personas, which affected their use behaviors. We present these findings, along with further analysis in the Appendix E.
5.2. RQ2: Comparison of User Engagement between Retrieval-based and Multi-persona Debate System
In this section, we analyzed participants’ eye movements within attitude-consistent vs. inconsistent AOIs to understand how they engaged with the two systems. We begin by analyzing the system differences then diving into how variations of eye movements between systems were influenced by two human factors, including topical familiarity and confirmation bias tendency (Section  5.2.1). Next, we present the analysis of the variance of eye movement across different behavioral groups (Section  5.2.2). As the eye movement data follows a normal distribution, our statistical analysis primarily applies parametric test.
5.2.1. Effects of bias mitigation across systems and human factors
Table  3 shows the analysis of ET metrics (i.e., total fixation counts and duration) between baseline and debate systems in how participants read the attitude-consistent vs. attitude-inconsistent information. Figure  4 shows total fixation duration on attitude-consistent vs. attitude-inconsistent information in both systems across different levels of topic familiarity and confirmation bias tendency. We introduce our statistical results as follows.
Interface
Fixation measure Attitude-consistent Attitude-inconsistent Paired t-test
Retrieval-based Duration 61073.67 67063.19 -1.324
Counts 277.69 307.39 -1.497
Multi-persona debate Duration 17642.06 20519.94 -2.229*
Counts 83.10 94.84 -2.421*
Table 3. Comparison of Interface Measurements for Attitude-consistent and Attitude-consistent Conditions. The significance levels are represented as p-value, 0 ¡ ‘***’ ¡ 0.001 ¡ ‘**’ ¡ 0.01 ¡ ‘*’ ¡ 0.05.
Figure 4. Comparison of total fixation duration on attitude-consistent and inconsistent information, displayed in box plots, between baseline and debate systems, analyzed across varying levels of topic familiarity and confirmation bias tendency levels.
System comparison. As shown in Table  3, our paired t-test revealed no significant difference in eye fixation counts and duration between attitude-consistent and attitude-inconsistent content in the baseline system, but a significant difference was observed with the debate system. Additionally, as both fixation counts and duration were statistically significant higher in the attitude-inconsistent information for the debate system, this indicates participants’ greater willingness to engage with diverse perspectives in the debate system than the baseline.
[…]
The ANOVA analysis conducted above on content-belief consistency but separately with topic familiarity and confirmation bias tendency reveal a consistent result: there is a clear difference on whether participants reading content consistent or inconsistent with their attitude when using the debate system, as reflected in their eye movements, but no difference found in the baseline. While topic familiarity yields no difference in both systems, confirmation bias tendency resulted in significant difference on the baseline but not the debate system. In combined with the system comparison results shown in the paired t-test (Table 3), we can conclude that the debate system effectively reduces selective exposure and increases user engagement with attitude-inconsistent content, while the baseline system activates confirmation bias tendencies, as evidenced by distinct eye movement patterns.
5.2.2. Effects of bias mitigation across user behavior patterns
Figure 5. Comparison of normalized fixation duration and fixation counts on attitude-consistent and inconsistent information across user behavior groups.
[…]
These results suggest that there is significant difference in the eye movements between behavior clusters within the debate system, As content-belief consistency did not significantly influence fixation counts, this suggests that user engagement levels were more closely tied to their behavior patterns when using the debate system rather than whether participants would read content associated with their prior beliefs. Additionally, as post-hoc analysis revealed that Critical Readers (group 2) had significantly higher fixation counts compared to the other groups, this suggests that Critical Readers were more visually engaged with the content. By associating with qualitative results, describing how Critical Readers actively engaged with various personas and generated content (Section  5.1), we can conclude that certain behavioral pattern, i.e., Critical Readers, show a clear sign of engagement demonstrated in their eye movement data.
5.3. RQ3: Comparison of User Belief Changes between Retrieval-based and Multi-persona Debate System
In this section, we first examine whether participant belief changed or remained unchanged across two systems and the influence of topic familiarity and confirmation bias tendency (Section 5.3.1, similar to our approach for RQ2). Afterward, we explore in greater depth in terms of bidirectional belief changes, i.e., whether participants became more or less supportive, because this helps suggest whether there is a shift toward increased or decreased polarization (Section 5.3.2). As the belief change data does not follow a normal distribution, we applies non-parametric test for analysis.
5.3.1. Effects of non-directional belief changes across systems and human factors
By visually presenting participant belief changes based on their topic familiarity (Figure 6, the two plots on the left) and confirmation bias tendency (Figure 6, the two plots on the right), we observed a clear trend indicating that participants with higher familiarity with the topic were less likely to change their beliefs. However, there is a less obvious trend shown in confirmation bias tendency. We conducted a generalized linear mixed-effects model analysis to quantify the statistical effects (model details were presented in Appendix F.2). The overall result demonstrates that there is a statistically significant difference across the various levels of topic familiarity and confirmation bias tendency on whether participants changed belief or remained unchanged, while the difference is not shown in interface conditions. We briefly describe the mixed effects for each variable in details as follows.
Figure 6. Percentage stacked belief changes bar charts between baseline vs. experiment across topic familiarity and confirmation biases tendency: there is a clear tendency showing that participants with higher topic familiarity were less likely to change their beliefs, as presented in both systems. However, no such trend was evident across different levels of confirmation bias.
System comparison: The system variable has an estimated positive coefficient of $0.2474$ ($SE=0.2565$, $z=0.964$ , $p=0.3349$). This suggests that participants using the debate system would slightly change their beliefs. However, this likelihood is very small and not statistically significant.
Topic familiarity: Topic familiarity has a statistically significant negative impact on belief change, with an estimated coefficient of $-0.6131$ ($SE=0.1369$, $z=-4.477$ , $p<0.001$). This suggests that as participant familiarity with a topic increases, they become significantly less likely to change their beliefs.
Confirmation bias tendency Confirmation bias tendency shows a statistically significant positive effect on belief change with an estimated coefficient of $0.4338$ ($SE=0.1873$, $z=2.315$ , $p<0.05$). This means that participants with higher levels of confirmation bias tendency are more likely change their beliefs, which is counterintuitive to our understanding. By examining the different proportions of belief changes across CB groups, we found that the low tendency group had the highest individual percentage of belief change (25.6%), but the combined belief change percentages of the medium and high tendency groups (19.1% + 20.6% = 39.7%) were greater than that. This cumulative effect may indicate that participant belief changes might not operate neatly into a three-level continuum of confirmation bias tendency.
Correlation among three variables: As parts of the generalized fixed-effect results, there is no moderate correlation among group interface, topic familiarity, and confirmation bias tendency, reflecting that these variables operated relatively independently in influencing belief change.
5.3.2. Effects of directional belief changes across systems and human factors
We examined how system conditions influence participant bi-directional belief changes, weighted by their confidence level, and explored this variance on topic familiarity and confirmation bias tendency.
Measurements Retrieval-based Multi-persona debate Wilcoxon
# — % Less supportive 19 — 11.8% 39 — 24.4% /
# — % More supportive 32 — 20.0% 21 — 13.1% /
# — % Unchanged 109 — 68.2% 100 — 62.5% /
M — Std Weighted changed 0.30 — 2.20 -0.34 — 2.72 1040.0*
Table 4. Bi-directional belief change between baseline and experiment condition: a larger number of participants became less supportive after using the experiment system compared to the baseline. The weighted change for the experiment was significantly lower than that of the baseline, suggesting that the experiment is more likely to lead participants to become less supportive of their initial stance. The significance levels are represented as p-value, 0 ¡ ‘***’ ¡ 0.001 ¡ ‘**’ ¡ 0.01 ¡ ‘*’ ¡ 0.05
Figure 7. Bidirectional belief changes box plots for the baseline and experimental interfaces across topic familiarity and confirmation biases tendency: Over both factors, while many participants did not alter their beliefs, which reduced the overall magnitude of changes in the box plots, for participants who did change their beliefs, the multi-persona debate system was more likely to lead participants to become less supportive than the baseline.
System comparison: By mapping bi-directional belief changes — less supportive, more supportive, and unchanged — on both systems (Table 4), we found that although many participants did not change their beliefs ($Retrieval_{unchanged}$ = $68.2\%$ , $Debate_{unchanged}$ = $62.5\%$), more participants became less supportive after using the debate than the baseline ($Debate_{less}$ = $24.4\%$ $Retrieval_{less}$ = $11.8\%$). Additionally, by conducting Wilcoxon sign-ranked test on the weighted belief change (in Table 4, the last row, where the mean value ¡0 indicates becoming less supportive and ¿0 indicates becoming more supportive), the weighted belief change for the debate was significantly lower than that of the baseline ($Debate_{weightedchange}$ = $-0.34$ $Retrieval_{weightedchange}$ = $0.30$, $X^{2}=1040.5$, $p<0.05$).
Topic familiarity: By visualizing participant weighted score for belief change across different levels of topic familiarity (Figure 7, the two plots on the left), we found many participants did not alter their beliefs, which reduced the overall magnitude of changes shown in the box plots. However, among participants who did change their beliefs, there is a lowered central tendency (median and the interquartile range) in the experiment compared to the baseline. This expands on earlier findings of significant difference across levels of topic familiarity (Section 5.3.1), demonstrating that, for participants who finally changed their belief, the debate system led them to become less supportive across different familiarity levels. Additionally, we also found the weighted belief change vary depending on different topics (additional analysis and findings are presented in Appendix F.3).
Confirmation bias tendency: By visually presenting participants’ bidirectional belief changes across different levels of confirmation bias tendency, we found a similar trend observed in the topic familiarity (Figure 7, the two plots on the right). While many participants did not change their beliefs, reducing the overall magnitude changes as presented in the box plot, for participants who did changed their beliefs, the debate system made user less supportive across three levels of confirmation bias tendency, whereas the baseline tended to increase it.
By combining the statistical results in both Section 5.3.1 and 5.3.2, we learned that while the system conditions did not yield a statistically significant difference on participant non-directional belief changes (i.e., whether participant belief changed or remained unchanged), for directional belief changes, there was a clear distinction. The debate system was more likely to cause participants to become less supportive of their initial stance, whereas the baseline tended to reinforce prior beliefs, making them more supportive. Additionally, user belief changes exhibited a statistically significant increase as topic familiarity decreased. While a statistically significant difference was also observed with confirmation bias tendencies, the trend did not fully align with our hypothesis. This discrepancy may be due to the three-level classification of bias tendency scores, indicating a need for further investigation in future studies.
6. Discussion
6.1. Enhancing Interaction
While our LLM debate system supports several forms of user interaction (e.g., topic selection, persona revision or creation, argument bookmarking, and controlling debate duration), participants offered several ideas to further increase interaction and engagement. Specifically, they sought greater ability to actively participate in and influence the debate.
For example, some participants suggested use of the search bar to allow them to inject their own opinions into an ongoing debate (P2-3). Similarly, some participants misunderstood the thumbs up and down bookmarking icons as allowing them to influence the ongoing debate and expressed a desire for this feature when realizing it was not supported (P26). Others requested the ability to highlight specific quotes from a persona content and prompt the persona to provide further information or supporting evidence (P25-26, 28, 39). A fact-checking feature was suggested to show the most recent evidence from the web for user-highlighted content, aiding in verification of arguments (P39).
6.2. Balancing Divergent Viewpoints and Convergent Solutions
While participants appreciated the system for broadening their understanding of controversial topics through multiple viewpoints, many expressed a desire for more concrete solutions to address the social issues being debated.
For example, some participants preferred concrete answers over the exchange of opinions. P13 noted personas“tend to stop at potential angles for the topic, but they don’t dive much deeper into more specific details.” P29 said “the chat didn’t address specific concerns and I couldn’t find a clear solution.” This led us to further reflect on when a system should offer diverse perspectives vs. more concrete answers. Providing convergent viewpoints can help spur decision-making but may also lead to automation bias in users too readily accepting system answers, something we sought specifically to avoid in our design. Bearing a more critical and constructive stance, we consider there are different design opportunities to balance divergent viewpoints with convergent solutions effectively.
First, convergent solutions could be presented when debate becomes congruent during the debate. For example, our participants noted that in the later rounds, the debate became less meaningful or began to repeat similar comments (P21, 26, 35, 40). Previous studies have reported similar findings, noting that LLMs tend to stop generating novel thoughts once they have established confidence in their responses — a phenomenon known as the “degeneration-of-thought” problem (Liang et al., 2023). While novel prompting techniques could be developed to help LLM continue generating new thoughts (Chen et al., 2023), or to allow participant inputs to illicit new generations, this might also be a sign for prompting convergent solutions instead of continued divergent brainstorming. Additionally, it may be necessary to distinguish between points of agreement and disagreement, and then generate solutions tailored to each (Shaikh et al., 2024a). Implementing this functionality could help transform the debate system for divergent thinking into more of a practical problem-solving tool. A practical feature would be automatically detecting when to wrap-up the debate and signal this to the user, improving our current design, which lacks guidance to the user when continued debate is unlikely to yield additional insights.
Second, while some contentious topics have little consensus (and so are very suitable for public debate), other topics have clearer scientific answers, such as whether climate change exists or the Earth is flat. In such cases, it can be not only misleading but harmful to present (and give standing) to arguments that continue to dispute well-accepted scientific facts. For example, fact-checking agencies often choose not to fact-check highly sensational claims for which the simple act of fact-checking would lend more credence and visibility to these claims than they merit (Phillips, 2018; Procter et al., 2023).
Additionally, integrated fact-checking support might not be needed if the system automatically showed the (detected) reliability of all statements presented, such as based on evidentiary support. This could enable users to better calibrate their assessment of arguments presented with low vs. strong reliability indicators. Alternatively, we could further constrain and filter the debate (personas and/or arguments) to only include arguments meeting a minimum reliability bar. In this case, the “debate” may present diverse personas and evidence supporting a given stance rather than arguing for different positions.
All that said, one can still imagine cases where fringe opinions might be very valuable to include. A science educator might want to know why flat-Earthers believe what they do so that their beliefs may be effectively countered. A scientist operating at the frontier of their field may wish to explore a new idea that challenges accepted knowledge (Liu et al., 2024). A writer may be seeking assistance in generating dialogue for a fictional story or script (Wan et al., 2024). As in so many things, context matters. Therefore, design should strive to flexibly support such different contexts and needs so that the debate includes appropriate, responsible, and accurate presentation of information that is fit for purpose.
6.3. Addressing Potential Bias in LLM-generated Personas
LLM simulation may exhibit many inaccuracies vs. reality due to hallucination, random errors, and consistent biases. Approximation of diversity in the real world may over- or under-representing certain groups and perspectives. Such bias often stems from the limitations of training data that incorporate existing social inequalities, stereotypes, or incomplete views of the population (Christian, 2020). Recent work has examined these misalignment across misinformation (Neumann et al., 2024), climate change (Santurkar et al., 2023), and politics (Taubenfeld et al., 2024).
In our study, we found that participants intentionally assessed whether LLM simulations accurately reflected the roles they represented (P19, 29, 31, 34, 38-39) and revised the descriptions if they considered simulations might be biased (P29, 34). Additionally, some participants even challenged the limits of LLM simulations by creating unconventional personas. For example, P34 created the persona of a foreigner with no ties to America to evaluate whether the LLM responses were both realistic and relevant to the federal corporate tax issue. P2 took on the persona of an animal to test if the LLM could effectively respond from an animal perspective on animal testing “I am surprised by the cat persona, it actually understands the description, hates dogs, and says things less logically.
Of course, users may hold their own biases toward certain social roles which could further reinforce their confirmation biases. The very flexibility our system enables users to customize the LLM personas and correct biases they find also means that they have the power to customize personas that reaffirm their existing beliefs. We cannot objectively evaluate how well participant customization reflects social roles within the demographic groups embodied by those personas. Future work should involve systematic socio-demographic studies to better understand, align, and compare user interpretations of personas with the biases present in LLM-generated personas (Kirk et al., 2024). Additionally, HCI research might consider how to balance giving users control over customization with ensuring representational diversity and fairness in LLM-generated personas.
6.4. Reducing Automation Bias in AI-Assisted Decision-Making
Although we did not explicitly assess the system’s impact on automation bias (Bertrand et al., 2022) (i.e., whether participants would automatically follow AI suggestions or engage in deliberate decision-making), the debate format of arguing opposing positions directly avoids providing simple answers to users that might be accepted without deliberate engagement. The debate format, with high customization, forces users to think critically and objectively assess the different arguments presented. This contrasts with the more passive role users typically take in traditional AI-assisted decision-making scenarios, where they might simply accept AI decisions (Buçinca et al., 2021; Lai et al., 2023).
Many participants intentionally structured a robust debate by selecting or customizing personas with different stances on the question (P14, 22, 24). P22 stated, “I felt like the descriptions that it gave would give a good debate” P24 added, “I’d probably change these personas a little more into what I feel like would really structure a good debate.” Although asking people to grapple with uncertainty may be uncomfortable, this encourages users to engage with multiple viewpoints and think critically to promote more deliberate decision-making. We were inspired in part by prior work using cognitive forcing functions to mitigate user over-reliance on AI decisions (Buçinca et al., 2021; Gajos and Mamykina, 2022). The debate system also offers this opportunity to help provide potential opposing viewpoints, fostering appropriate reliance in AI-assisted group decision-making (Chiang et al., 2024).
In general, participant comments suggested that they found it engaging and enjoyable to engage in this way (P12, 24, 32, 38), reflecting on their prior stances on social issues and becoming aware of their own biases (P33). P38 remarked, “It was interesting and nuanced to control an LLM, something I’ve never experienced before.” P33 noted, “I think I still have some personal bias in terms of what I look for more.” We aim to combine increasing exposure to diverse views — with an interface that promotes this rather than ignoring perspectives users might dislike — and avoiding spoon-fed answers that can be accepted without thought. By doing so, we hope to foster critical thinking and reduce reliance on automation, ultimately enhancing the decision-making process.
6.5. Limitations
The discussion above highlights several key limitations of our study. In this section, we outline these limitations concerning system design and development (Section 6.5.1), sampling and test time (Section 6.5.2).
6.5.1. System design and development
As described in Section 6.1, our current multi-persona debate system lacks some interactive features to increase user engagement, such as the ability to nudge debate conversations by incorporating user feedback and new inputs. Additionally, system improvements are needed to enhance debate quality (Section 6.2), such as integrating RAG to reduce potential hallucination problems, refining prompting techniques to minimize repetitive responses from personas, and facilitating a natural closure in the debate (e.g., “I have nothing new to add”).
Additionally, because the system has not been evaluated for potential biases in its representations across topics and the persona, it is highly likely that biases in the LLM could impact the results of our user testing. Thus, in terms of system design, future studies could incorporate these interactive features to enhance user engagement and integrate more robust techniques to address potential biases within the system.
6.5.2. Sampling and test time
We also recognize that our findings are limited by the participant pool, which primarily consisted of college students. This focus restricts the generalizability of our results to a broader population. College students often lack real-world experience with certain controversial topics, which may affect their level of engagement and reduce the diversity of perspectives represented in the study. Expanding participant diversity in future research could not only broaden the range of perspectives for evaluating system effectiveness, enhancing generalizability, but also inspire new usage behaviors or scenarios for the debate system.
Additionally, our study restricted participant task time in the lab, which may have limited their ability to fully explore long-term interaction patterns with the system. Extended use of the system could potentially reveal different engagement behaviors and belief changes that were not captured within the current study time constraints. Therefore, both sampling and time limitations suggest that future studies should consider diverse samples and longer engagement periods to capture a more comprehensive understanding of the debate interface impact across varied user groups.
7. Conclusion
This study introduces and evaluates an LLM-powered multi-persona debate system designed to mitigate confirmation bias in information access.
The system presents diverse viewpoints through an interactive, persona-based interface where users engage with simulated debates across multiple perspectives on controversial topics. Through an in-depth, within-subject, mixed-method user study, we demonstrate that the multi-persona debate system significantly reduces confirmation bias by encouraging users to invest more attention in content that challenges their initial beliefs, as evidenced by the eye-tracking data and behavioral measures.
Our results provide valuable insights for designing future LLM-based systems to foster open-minded engagement with diverse viewpoints, encouraging unbiased information processing on complex or controversial topics.
8. Acknowledgments
This research is funded in part by Microsoft and by Good Systems [^10] https://goodsystems.utexas.edu, a UT Austin Grand Challenge dedicated to developing responsible AI technologies, as well as by the UT Austin School of Information. We extend our special thanks to Gauri Kambhatla for her assistance regarding polarization research, and to Anubrata Das and Sooyong Lee for their valuable feedback and insights. We are also grateful to our participants, without whom this research would not have been possible. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
Appendix
Appendix A Prompts
A.1. Prompt Placeholders
Throughout our prompts, we use placeholders, which are automatically replaced by their corresponding value as defined in our initial configuration. For all our current version of the system these values are:
  1. (1)
    [TOPIC]: The user-input topic to focus persona generation and debate discussion around.
  2. (2)
    [NUM_PERSONAS]: Initial requested number of personas
  3. (3)
    [CURRENT_PERSONAS]: Current personas in JSON format
  4. (4)
    [NAME]: Name of the persona
  5. (5)
    [DESC]: Description of the persona
  6. (6)
    [HISTORY]: Transcript of the debate in JSON format
  7. (7)
    [LIMITER]: Sentence telling the LLM how long their response should be (ie. Limit your response to 5 sentences).
A.2. Persona Generation
A.2.1. Generating an initial set of personas
Given the topic, [TOPIC], create a roundtable debate of different personas to show key perspectives of it. Output the personas as a list of JSON objects. Each JSON object should have the following structure:
{ "title": <title of the Persona>,
"description": <Brief Description of the Persona, only describing their background (rather than their stance)>,
"emoji": <Single emoji representation of the perspective the persona represents>}.
Ensure that the output is formatted as a valid JSON.
Please generate exactly [NUM_PERSONAS] personas.
A.2.2. Adding an Additional Persona
Given the topic, [TOPIC], we are creating a roundtable debate of different personas to show key perspectives of it.
Currently, we have personas as follows: [CURRENT_PERSONAS] Please output exactly one additional persona. Your output should have the following JSON structure:
{"title": <title of the Persona>,
"description": <Brief Description of the Persona, only describing their background (rather than their stance)>,
"emoji": <Single emoji representation of the perspective the persona represents>}.
Ensure that the output is formatted as a valid JSON.
A.3. Persona Response
A.3.1. Starting a debate
You are in an roundtable debate on the topic [TOPIC].
You are [NAME], who is [DESC].
Please start the debate by concisely presenting your argument for your stance on the topic. [LIMITER]
A.3.2. Continuing the debate
You are in a continuing roundtable debate on the topic [TOPIC].
You are [NAME], who is [DESC].
Here is the transcript of the debate so far  [HISTORY]
Please continue to debate the others, concisely supporting your stance on the topic. [LIMITER]
Appendix B Measurements for data collection
B.1. Pre task questionnaire
  1. (1)
    Please indicate your current stance on the given topic? [Supportive; Neutral/ Undecided; Opposed]
  2. (2)
    How confidence are you for your stance and reason? [Very low; Slightly low; Moderate; Slightly high; Very high]
  3. (3)
    How familiar are you with the topic? [Not familiar; Slightly familiar; Moderately familiar; Familiar; Very familiar]
B.2. Post task questionnaire
  1. (1)
    Please indicate your current stance on the given topic? [Supportive; Neutral/ Undecided; Opposed]
  2. (2)
    How confidence are you for your stance and reason? [Very low; Slightly low; Moderate; Slightly high; Very high]
B.3. Interview questions
  • When you were using the retrieval-based system,
    1. (1)
      what kind of information interested you most?
    2. (2)
      What document did you decide to read?
    3. (3)
      What were you looking for in the document page?
    4. (4)
      Which interface did you find most useful?
    5. (5)
      Did you find the content trustful/ useful? Why?
  • When you were using the multi-persona debate system,
    1. (1)
      How did you decide which persona to select to engage in the debate?
    2. (2)
      How did you decide how to switch agents between debate rounds?
    3. (3)
      Why did you change (not change) the personas between rounds?
    4. (4)
      How did you decide which persona to change/ keep between rounds?
    5. (5)
      How many rounds of debate did you usually read? Why did you want to proceed to the next round? When would you stop generating next round?
    6. (6)
      What kind of content interested you most/ helped you to explore the topic and make decision?
    7. (7)
      Did you find the bookmark function useful? Why?
    8. (8)
      How did you decide which information to save in the bookmark?
    9. (9)
      Did you find the LLM-generated information useful and trustful?
  • In general, how did you use these two systems to explore the diverse perspectives for the given topics? Where did you found most useful for your exploration?
  • For the two systems you used in this study,
    1. (1)
      Were you satisfied with the experience when using the systems?
    2. (2)
      Were you overloaded by information?
    3. (3)
      Was the design effective?
    4. (4)
      which one do you prefer?
  • Do you think your prior knowledge would influence how you use the systems?
    1. (1)
      When you were familiar with the topic, how would you use the systems?
    2. (2)
      when you were unfamiliar with the topic, how would you use the systems?
  • When exploring with the system, what factors/ information would make you shift your perspective?
Appendix C Baseline Retrieval-based System
Figure 8. Screenshot of Daxenberger et al. (2020)’s “ArgumentSearch ” system. When a user enters a query, such as “Do violent video games contribute to youth violence?” the system identifies relevant articles and presents content snippets representing both sides, including those in favor (PRO) and those opposed (CON) to the topic.
ArgumentSearch is a retrieval-based search engine that identifies articles from the open web and organizes content snippets according to their stance on the user query. We chose ArgumentSearch over standard search like Google or conversational search tools like Perplexity because it specifically organizes articles by stance (supportive or opposed), aligning with the core idea of presenting content from both sides of an issue to mitigate confirmation biases (as described in Section 2.2). By comparing this system with the multi-persona debate system, we can more effectively evaluate the mere impact of personas while mitigating the common design effect of one-dimensional ranking in search results, which tends to naturally cause a reinforcement of user confirmation bias and echo chambers during the traditional information-seeking process.
Appendix D Participant Demographics
ID Gender Age Highest Education Discipline
1 Female 23 Bachelor’s degree Applied sciences
2 Female 22 High school or equivalent Social sciences
3 Female 24 Bachelor’s degree Engineering
4 Male 28 Bachelor’s degree Natural sciences
5 Female 19 High school or equivalent Social sciences
6 Female 24 Bachelor’s degree Social sciences
7 Female 28 Bachelor’s degree Other
8 Female 20 High school or equivalent Natural sciences
9 Male 20 High school or equivalent Natural sciences
10 Female 29 Bachelor’s degree Social sciences
11 Male 25 Master’s degree Applied sciences
12 Female 25 Master’s degree Natural sciences
13 Male 32 Master’s degree Natural sciences
14 Female 18 High school or equivalent Social sciences
15 Female 29 Professional degree (MD, JD, etc.) Natural sciences
16 Male 18 High school or equivalent Natural sciences
17 Male 29 Doctoral degree Natural sciences
18 Non-binary/third gender 20 High school or equivalent Natural sciences
19 Male 18 High school or equivalent Applied sciences
20 Male 19 Bachelor’s degree Engineering
21 Female 25 Bachelor’s degree Engineering
22 Non-binary/third gender 20 High school or equivalent Natural sciences
23 Male 26 Master’s degree Information Science
24 Male 19 High school or equivalent Engineering
25 Female 19 High school or equivalent Applied sciences
26 Male 18 High school or equivalent Engineering (B.S.)
27 Male 19 High school or equivalent Natural sciences
28 Female 20 High school or equivalent Advertising
29 Male 20 High school or equivalent Applied sciences
30 Female 34 Master’s degree Humanities
31 Female 20 High school or equivalent Social sciences
32 Male 20 High school or equivalent Engineering
33 Male 21 High school or equivalent Applied sciences
34 Male 19 High school or equivalent Applied sciences
35 Female 25 Master’s degree Natural sciences
36 Female 26 Master’s degree Information Studies
37 Non-binary/third gender 20 High school or equivalent Engineering
38 Male 20 High school or equivalent Applied sciences
39 Male 32 Bachelor’s degree Computer Science
40 Female 20 High school or equivalent Applied sciences
Table 5. Participant information
Appendix E Additional Findings regarding User Behavior Patterns
Beyond the main behavior clusters outlined in Section 5.1, in this section, we provide additional analysis, describing the topical differences within these groups (Section E.1) and additional qualitative findings regarding how our participants select and revise personas (Section E.2).
E.1. Topical differences among user behavioral clusters
Figure 9. Distribution of user behavior patterns across different topics.
Additionally, when we separated these behavioral groups across different topics, we observed significant differences. As shown in Figure 9, for several topics, such as “Is vaping with E-cigarettes safe?” “Should the Federal corporate income tax rate be raised?” “Should TikTok be banned in school?” and “Do violent video games contribute to youth violence?”, the distribution of all three groups share a similar pattern [^11] The distribution of behavior patterns in the topic of “Should animals be used for scientific or commercial testing?” is similar to that of previous topics, but with a variation: there is a higher number Multi-round explorers compared to Critical readers. Among these topics, the Critical readers emerged as the majority group, followed by a smaller number of Multi-round explorers, and then Active observers. For other topics, including “Should humans colonize space?” “Should abortion be legal?”, and “Should the United States implement a universal basic income?”, the Multi-round explorers were overwhelmingly the dominant group, with only a few participants classified as Active observers or Multi-round explorers. It is important to note that there was a significant shift of participants from Critical readers to Multi-round explorers within these two topic groups. This change indicates that participants began allocating more of their time to generating new rounds of debate and relatively less time to reading and analyzing the content. This shift may have happened because participants sought more in-depth or novel insights on these specific topics, encouraging them to generate more rounds of debate. Overall, these results suggest that topic-specific factors do influence participant engagement patterns in using the debate system.
E.2. Persona expertise, individuality, and diversity in selecting and revising personas
From the post-task interviews, we found three additional factors that participants frequently mentioned in selecting and revising personas, including persona expertise, individuality, and the overall diversity of personas when selecting and editing them to construct the debate. These factors were closely tied to how participants conceptualized the controversial topics being discussed.
First, many participants preferred selecting both expert and individual personas. Expert personas were more trusted to provide factual information, evidence, and other logical and reliable arguments to the controversial topic (P11, 14-15, 22, 24, 33, 36, 38-39). P39 explained that for the topic of “TikTok ban,” experts could provide opinions rather than emotional responses “Parents who are concerned about their kid is great, but I don’t expect them to have all of the data necessary to make an informed decision. I thought there would be experts discussing things logically and providing evidence-based information, as opposed to information driven by emotions.” In contrast, individual personas could provide direct and authentic arguments (P11, 22, 24-25, 33, 38). For example, P38 said “Whether TikTok should be banned from school. I really want from people who are in the school specifically because it’s important to see the people who are actually being affected to know their pain points. So that’s why I picked the teacher, the student, and the administrator.
Additionally, as participants were aware that potential bias could underlie different personas, participants aimed to create a balanced representation of opinions by selecting personas that supported different sides of the topic. On one hand, this approach could be an effective strategy for generating counterarguments (P14, 24, 26, 28, 33, 35-36), as P36 explained, “I tried to balance the number of people who supported the topic with those who opposed it, to sort of weigh which perspective truly aligns with my own thinking and against it.” On the other hand, some would deliberately balance opinions that might be disproportionately represented on one side in real-world discussions (P22, 24, 33). P33 pointed out that“If I’ve been hearing a lot of one point of view, I’ll try to keep generating a persona until I find one that opposes it.
Appendix F Additional Statistical Results
F.1. Descriptive user behavior data in debate system
Mean Std. Deviation Minimum Maximum
Debate rounds $1.865$ $0.843$ $1.000$ $5.000$
Persona selected (pro) $1.734$ $1.195$ $0.000$ $6.000$
Persona selected (con) $1.973$ $1.300$ $0.000$ $9.000$
Debate duration $172.847$ $75.540$ $58.254$ $352.763$
New persona $0.974$ $1.081$ $0.000$ $5.000$
Persona edits $0.187$ $0.433$ $0.000$ $3.000$
Bookmark visits $0.121$ $0.274$ $0.000$ $1.000$
Bookmark made $2.628$ $2.689$ $0.000$ $9.000$
Table 6. Descriptive statistics of interactive behavior (mentioned in Section  4.4) in the multi-persona debate system.
F.2. Generalized linear mixed-effects model analysis
Variables Estimate Std. Error Z value P value
(Intercept) -0.3109 0.6324 -0.492 0.6230
UI 0.2474 0.2565 0.964 0.3349
TF -0.6131 0.1369 -4.477 ***
CB 0.4336 0.1873 2.315 *
Table 7. Generalized linear fixed-effect results: both topic familiarity (TF) and confirmation bias tendency (CB) has a statistically significant impact on belief change. The significance levels are represented as p-value, 0 ¡ ‘***’ ¡ 0.001 ¡ ‘**’ ¡ 0.01 ¡ ‘*’ ¡ 0.05.
Variables (Intercept) UI TF
UI -0.633 / /
TF -0.404 0.004 /
CB -0.469 0.021 -0.260
Table 8. Correlation among three variables and belief change: All three variables show a moderately negative relationship with the intercept while each variable operated relatively independently in influencing belief change.
We conducted a generalized linear mixed-effects model to analyze whether participant belief changed or remained unchanged varies on participant topic familiarity and confirmation bias tendency across two systems (Section 5.3.1). In this model, belief changes were treated as a binomial outcome (changed vs. unchanged), with user and topic included as random effects. The fixed effects included user confirmation bias tendency (CB), topic familiarity (TF), and the user interface (UI). We used the glmer function of the R package lme4121212https://cran.r-project.org/web/packages/lme4/index.html with the formula: $BeliefChange\simeq CB+TF+UI+(1|user)+(1|topic)$. We code higher value for each variable: a higher score on CB and TF reflects increased tendency on confirmation bias and topic familiarity; the experiment condition is coded with a higher value than the baseline.
F.3. Distribution of bi-directional belief changes across topics
ID Topics Groups Bi-directional belief changes (# — %)
less supportive more supportive Unchanged
1 Should animals be used for scientific or commercial testing? Retrieval 2 — 10.0% 6 — 30.0% 12 — 60.0%
Debate 3 — 15.0% 2 — 10.0% 15 — 75.0%
2 Should humans colonize space? Retrieval 2 — 10.0% 7 — 35.0% 11 — 55.0%
Debate 5 — 25.0% 4 — 20.0% 10 — 55.0%
3 Is vaping with E-cigarettes safe? Retrieval 2 — 10.0% 1 — 5.0% 17 — 85.0%
Debate 4 — 20.0% 0 — 0.0% 16 — 80.0%
4 Should abortion be legal? Retrieval 0 — 0.0% 1 — 5.0% 19 — 95.0%
Debate 0 — 0.0% 3 — 15.0% 17 — 85.0%
5 Should the Federal corporate income tax rate be raised? Retrieval 4 — 20.0% 3 — 15.0% 13 — 65.0%
Debate 13 — 65.0% 2 — 10.0% 5 — 25.0%
6 Should the United States implement a universal basic income? Retrieval 2 — 10.0% 8 — 40.0% 10 — 50.0%
Debate 1 — 5.0% 6 — 30.0% 13 — 65.0%
7 Should TikTok be banned in school? Retrieval 4 — 20.0% 3 — 15.0% 13 — 65.0%
Debate 2 — 10.0% 4 — 20.0% 14 — 70.0%
8 Do violent video games contribute to youth violence? Retrieval 3 — 15.0% 3 — 15.0% 14 — 70.0%
Debate 11 — 55.0% 0 — 0.0% 9 — 45.0%
Table 9. Bi-directional belief changes across different topics: The retrieval-based and multi-persona systems have varying effects on user belief changes depending on the topic. For topics 1 and 2, the retrieval-based system tends to reinforce participants’ previous stances, making them more supportive (highlighted as orange). In contrast, for topics 5 and 8, the multi-persona debate system leads participants to become less supportive of their initial stance (highlighted as lightblue). Both systems only show subtle impacts on other topics, but they both lead to participants becoming more supportive on topic 6 (highlighted as gray).
By mapping participant bidirectional belief changes across different topics (Table 9), we found that the yielded system effects on weighted belief changes also vary on the topic. For example, in topics “Should animals be used for scientific or commercial testing?” and “Should humans colonize space?” (highlighted as orange) participants became more supportive of their previous stances after using the baseline. In contrast, for topics “Should the Federal corporate income tax rate be raised?” and “Do violent video games contribute to youth violence?” (highlighted as blue) participants became less supportive of their previous stance after using the experiment. This suggests that user belief changes differ depending on the topic, with some topics being particularly sensitive.

### v3

@@
 1. Introduction
-Large language models (LLMs) offer a novel design space to explore new ways to support human information seeking.
+Large language models (LLMs) offer a novel design space to explore new ways to support human information-seeking.
 For instance, commercial search engines and services like ChatGPT and Perplexity can synthesize information from multiple sources to provide direct answers to user questions [^1] ChatGPT: https://openai.com/index/chatgpt/, Perplexity: https://www.perplexity.ai/..
 Recent work has also explored leveraging the internal knowledge and reasoning capabilities of LLMs to simulate human personas with diverse perspectives (Prpa et al., 2024; Park et al., 2022; Chen et al., 2024; Khan et al., 2024).
-Imagine a user could generate a set of LLM personas on-demand to debate any topic of interest from multiple perspectives.
-How should we design such a system to maximize potential benefit to information seekers (e.g., information diversity) while also minimizing associated risks (e.g., LLM hallucination and biases interacting with human biases)?
-In traditional information-seeking process, users retrieve facts, evidence, and other information from various sources, often guided by their pre-existing beliefs.
+Imagine a user could generate a set of LLM personas on demand to debate any topic of interest from multiple perspectives.
+How should we design such a system to maximize potential benefits to information seekers (e.g., information diversity) while also minimizing associated risks (e.g., LLM hallucination and biases interacting with human biases)?
+In the traditional information-seeking process, users retrieve opinions, evidence, and other information from various sources, often guided by their pre-existing beliefs.
 This can lead to confirmation bias (Berthet, 2021), where users favor information that supports their prior assumptions.
-Such biases are prevalent not only in standard information access systems, like search engines, but also in conversational interactions, where users actively shape the direction of the discourse (Sharma et al., 2024; Khan et al., 2024).
-However, LLM-powered diverse information-seeking could be intentionally designed to provide broad coverage and present contrasting views, regardless of prior user beliefs.
-Recent design explorations of multi-persona systems have been encouraging, with subjective or indirect user feedback suggesting mitigated confirmation bias (Park et al., 2023a; Chiang et al., 2024; Zhang et al., 2024).
-In this work, we explore a novel information seeking experience: letting users launch a debate on any topic of interest in which generated LLM personas present diverse perspectives on the topic.
-We explore the active and autonomous nature of LLM agents in this setting as they present information and arguments, creating an argumentative experience between humans and AI.
-In focusing the user experience on exposure to diverse perspectives, we particularly investigated whether exposure to diverse information via generated personas could help reduce user confirmation bias.
-Participants in our study launched LLM debates on controversial topics spanning social issues (e.g., ”Should abortion bans be illegal?”) to imaginary scenarios (e.g., ”Should space be colonized?”) We first assessed initial user attitudes toward these topics, then asked users to engage with a baseline search system and our multi-persona debate system.
-We used a mixed-methods approach to capture participants’ cognitive perceptions, such as real-time eye-tracking metrics, behavioral data, such as user interactions, and qualitative feedback through retrospective think-aloud sessions.
-These data not only help assess the relative performance of our multi-persona debate system, but also address limitations of other recent exploratory studies on LLM-persona systems, which primarily relied on self-reported metrics and subjective feedback (Park et al., 2023a; Zhang et al., 2024).
-Our study thus complements this prior work in moving toward a more comprehensive view of how use of LLM personas can impact user information consumption.
-Our study investigates three key research questions:
-RQ1:
-What is the overall user experience when conceptualizing controversial topics through a multi-persona debate interaction simulated by LLM?
-RQ2:
-Compared to the retrieval-based system, how effective is the multi-persona debate system in mitigating confirmation bias?
-RQ3:
-Compared to the retrieval-based system, how does the multi-persona debate system finally influence user pre-existing beliefs?
+Such biases are prevalent not only in standard search engines like Google, but also in newer conversational search systems powered by LLMs, where users can actively steer the discourse, reinforcing the exposure of information that favors their beliefs (Sharma et al., 2024; Khan et al., 2024).
+In another line of work (Park et al., 2023a; Chiang et al., 2024; Zhang et al., 2024), scholars explore LLM-enabled multi-persona information-seeking, which presents contrasting viewpoints by simulating different stakeholder perspectives, preventing users from unintentionally seeking evidence that validates their prior beliefs.
+This design promotes broader and more balanced information coverage with the potential to address confirmation bias as presented in the traditional and newer search systems.
+In our study, we further explore this emergent information-seeking experience to understand how it directly reduces confirmation biases.
+We argue that the user experience offered by multi-personas is particularly argumentative, as simulated personas provide users a constructive debate experience on any topic of interest, presenting diverse perspectives.
+While recent work has demonstrated the potential of using multi-personas to reduce echo chambers (Zhang et al., 2024) and to encourage critical thinking in group decision-making (Chiang et al., 2024), these studies primarily rely on subjective measurements, such as verbal accounts and user ratings, to assess effectiveness.
+Although these self-reports provide valuable user insights, they offer limited cognitive evidence for confirming the actual reduction of confirmation bias.
+To address this, we integrate eye-tracking metrics — supported by established correlations between visual attentional patterns and confirmation biases (detailed in Section 2.2) — as an objective measure to bridge subjective feedback and behavioral cognition.
+Participants in our study launched LLM-generated multi-persona debates on controversial topics spanning social issues (e.g., “Should abortion bans be illegal?”) to imaginary scenarios (e.g., “Should space be colonized?”).
+We first assessed participant attitudes toward eight controversial social topics and then asked them to use a baseline search system and the multi-persona debate system.
+We used a mixed-methods approach to capture their cognitive perceptions (i.e., real-time eye-tracking metrics), behavioral data (i.e., user interactions), and qualitative feedback through interviews and retrospective think-aloud.
+These data not only help assess the relative performance of the multi-persona debate system but also address limitations of other recent exploratory studies on LLM-persona systems, which primarily relied on self-reported metrics and subjective feedback (Park et al., 2023a; Zhang et al., 2024).
+Our study complements this prior work in moving toward a more comprehensive view of how the use of multi-personas can impact user information consumption.
+We investigate three key research questions: RQ1: What is the overall user experience when conceptualizing controversial topics through a multi-persona debate interaction simulated by LLM?
+RQ2: Compared to the retrieval-based system, how effective is the multi-persona debate system in mitigating confirmation bias?
+RQ3: Compared to the retrieval-based system, how does the multi-persona debate system finally influence users pre-existing beliefs?
 Through a combined analysis of both quantitative and qualitative data, our findings reveal that the argumentative experience powered by LLM personas helps foster constructive deliberation on contrasting viewpoints, making individuals more likely to consume content that challenges their pre-existing beliefs.
 This happens because persona-based debates present information in a more dynamic, appealing, and persuasive manner, reducing user cognitive load and enhancing interactivity so that people are willing to engage with challenging content.
-In summary, our work investigates persona-based debate systems to gain deeper insights into human cognitive perceptions and behavioral patterns when interacting with them.
+In summary, our work investigates multi-persona debate systems to gain deeper insights into human cognitive perceptions and behavioral patterns when interacting with them.
 We assess the effectiveness of these systems in reducing confirmation bias during information seeking and offer broader design implications to guide the future development of persona-based systems.
@@
 2.1. The Social Simulation and Role-Play of LLM
-Over the past five years, LLMs (e.g., GPT, Llama, Gemini, etc.) have rapidly advanced and today represent the state-of-the-art in performance across various natural language tasks [^2] GPT: https://openai.com/index/gpt-4/, Llama: https://www.llama.com/, Gemini: https://gemini.google.com/.
-Additionally, by being trained on the increasing amounts of web data and computational resources (Kaplan et al., 2020), LLMs have developed strong reasoning and association abilities, as well as internal knowledge that captures social knowledge and cultural norms.
-To instruct and tailor content generated by LLMs, prompting is a specific technique used by practitioners, providing initial inputs or cues to a language model (Liu et al., 2023).
-Researchers have explored creative ways to assess and extend the human-like capabilities of LLMs through prompting (Wang et al., 2021; Wei et al., 2022; Yao et al., 2022).
-One novel application is the development of agent-based systems where LLMs are used to create highly autonomous agents.
-LLM personas (or Social Simulacra) are a prominent example.
-Specifically, by incorporating different beliefs, desires, goals, and characteristics into the prompts, LLMs can role-play to simulate human reactions and behaviors (Shanahan et al., 2023; Park et al., 2022; Chen et al., 2024; Tseng et al., 2024).
-This approach offers new opportunities for enhancing user experiences in Computer Supported Cooperative Work (CSCW) and other areas of Human-Computer Interaction (HCI).
-In these fields, generative agents can provide more dynamic, interactive, and autonomous user experiences (Park et al., 2023b).
-Researchers have explored different use cases, such as using social simulacra to conduct user studies (Argyle et al., 2022; Park et al., 2022), creating generative agents in an interactive sandbox environment (Park et al., 2023b), or directly using LLMs as judges to assist decision-making tasks (Pan et al., 2024), ranging from data annotation (Tan et al., 2024) to language evaluation (Wang et al., 2023).
-However, researchers also highlight important challenges in the real-world experiments that employ LLM as the core technique to design persona systems, as well as agent-based systems in general.
-For example, in particular on the development of social simulacra, LLMs can reflect polarized human opinions regarding different social issues, including misinformation (Neumann et al., 2024), climate change (Santurkar et al., 2023), and politics (Taubenfeld et al., 2024).
+LLMs (e.g., GPT, Llama, Gemini, etc.) have rapidly advanced and today represent the state-of-the-art in performance across various natural language tasks [^2] GPT: https://openai.com/index/gpt-4/, Llama: https://www.llama.com/, Gemini: https://gemini.google.com/.
+By being trained on the increasing amounts of web data and computational resources (Kaplan et al., 2020), these models have developed strong reasoning and association abilities, as well as internal knowledge that captures social knowledge and cultural norms.
+To instruct and tailor content generated by LLMs, practitioners commonly use prompting by providing initial inputs or cues to steer the generation (Liu et al., 2023).
+Researchers have proposed various methods to enhance LLMs’ capabilities through prompting techniques, such as chain-of-thought or interactive reasoning (Wang et al., 2021; Wei et al., 2022; Yao et al., 2022).
+A notable application in this domain is the creation of agent-based systems.
+LLM-enabled autonomous agents are designed to simulate complex social behaviors.
+A prominent example is LLM personas (or Social Simulacra), which enable dynamic, human-like interactions within simulated environments.
+In earlier papers that use LLM as social simulacra, Park et al.
+(2022) embedded traditional UX personas into the prompt, which include “interests (e.g., hobbies, jobs) and personality traits (e.g., kind, bully)”.
+Their later work (Park et al., 2023b) extends to the incorporation of contextual “summary description (e.g., name, traits, and recent experiences)”.
+Recent work follows this trend.
+For example, Park et al.
+(2023a) use purchase criteria to create assistant agents for camera shopping, Zhang et al.
+(2024) incorporate character traits “(gender, age, occupation, and education)” to create virtual conversational agents.
+In related survey paper, Chen et al.
+(2024) synthesize that persona prompts are dynamically constructed by merging static sociodemographic traits (e.g., age, profession) with emergent behavioral traits shaped through iterative interactions with simulated environments.
+Additionally, debate is a key interaction that characterizes how LLM multi-personas engage with one another.
+This implementation also varies.
+For example, debates can be structured as simple opposition to user decisions (Chiang et al., 2024), multi-option decision-making scenarios where personas negotiate choices (Park et al., 2023a), or nuanced conversational exchanges based on sociodemographic differences in the persona configuration (Zhang et al., 2024).
+The creation of LLM personas offers new opportunities for enhancing user experiences Researchers have explored different use cases, such as using social simulacra to conduct user studies (Argyle et al., 2022; Park et al., 2022), creating generative agents in an interactive sandbox environment (Park et al., 2023b), or directly using LLMs as judges to assist decision-making tasks (Pan et al., 2024), ranging from data annotation (Tan et al., 2024) to language evaluation (Wang et al., 2023).
+In particular, the debate of LLM multi-personas could be useful in mitigating online polarization and mis/disinformation, one of the core research areas within Computer Supported Cooperative Work & Social Computing (CSCW).
+We elaborate on this in Section 2.3.
+Researchers also highlight challenges in real-world experiments that employ LLM as the core technique to design persona or agent-based systems.
+For example, on the development of social simulacra, LLMs can reflect polarized human opinions regarding different social issues, including misinformation (Neumann et al., 2024), climate change (Santurkar et al., 2023), and politics (Taubenfeld et al., 2024).
+Additionally, the potential misuse of LLMs could further exacerbate these issues, amplifying existing problems rather than mitigating them (Zhou et al., 2023; Bender et al., 2021).
 More importantly, the broad idea of agent-based systems also raises important ethical questions about the idea of replacing humans (Prpa et al., 2024).
@@
 Ongoing research is needed to investigate both the advantages and challenges of applying these systems to different real-world scenarios.
-2.2. Confirmation Biases in Information Seeking
+2.2. Confirmation Biases and Debiasing Strategies in Information Seeking
+2.2.1. The Definition and Measurements of User Confirmation Biases
 Confirmation Bias is one of the cognitive biases (Berthet, 2021) which refers to the human tendency to seek and interpret evidence that supports their existing beliefs, expectations, or hypothesis (Nickerson, 1998), regardless of the plausibility of those beliefs (Edwards and Smith, 1996).
-Confirmation bias consists of two aspects: the cognitive aspect (disposition) and the motivational aspect (manifestation) (Berthet et al., 2022).
-The cognitive aspect of confirmation bias focuses on human tendencies (Berthet et al., 2022; Baron, 2023).
-It represents a stable psychological disposition that reflects an individual’s tendency toward confirmatory information seeking and evaluation, which can be reliably measured through validated psychometric instruments (Ardi and Pradiri, 2021; Rassin, 2008; Berthet, 2021).
+It consists of two aspects: the Cognitive aspect and the Motivational aspect (Berthet et al., 2022).
+The cognitive aspect focuses on human tendencies (Berthet et al., 2022; Baron, 2023).
+It represents a stable psychological disposition that reflects an individual tendency toward confirmatory information seeking and evaluation, which can be reliably measured through validated psychometric instruments (Ardi and Pradiri, 2021; Rassin, 2008; Berthet, 2021).
 The motivational aspect focuses on the relationship with users’ prior beliefs (Baron, 1995).
-This aspect relates to users’ pre-existing beliefs within specific contexts and manifests dynamically during information-seeking activities, where individuals tend to select information aligning with their attitudes, allocate more attention to confirming content, and assign greater weight to belief-consistent evidence during evaluation (Vedejová and Čavojová, 2022).
-Previous research has shown that confirmation bias is widely manifested in seeking behavior, influencing how users interact with information.
+This aspect contextualizes user beliefs within specific contexts and manifests dynamically during information-seeking activities, such as individuals tend to select information aligning with their attitudes, allocate more attention to confirming content, or assign greater weight to attitude-consistent evidence during evaluation (Vedejová and Čavojová, 2022).
+Our study measurements cover both aspects to evaluate the reduction of confirmation bias (see details for RQ2 and RQ3 in Section 4.4).
+Previous research has explored the influence of confirmation bias on how users seek, browse, select, and engage with attitude-consistent information versus attitude-inconsistent information.
+In this context, confirmation bias was quantified by comparing either the number or proportion of selected attitude-consistent versus attitude-inconsistent items (Betsch et al., 2001; Jonas et al., 2008; Schulz-Hardt et al., 2000; Ling et al., 2017).
+For example, Rieger et al.
+(2021a) calculated the ratio of attitude-consistent to inconsistent selections.
+Huang et al.
+(2012a) classified users as demonstrating confirmation bias if more than 50% of their selected items aligned with their prior beliefs.
+In addition to methods relying on self-reporting and interaction logs, more unobtrusive methods, such as eye-tracking, were also used in measuring confirmation bias.
+Eye-tracking is a well-established technique in information retrieval (IR) research (Mostafa and Gwizdka, 2016).
+It relies on the eye-mind link hypothesis — visual attention follows eye movements (Just and Carpenter, 1980), such as record fixations, scanpaths, pupil dilation, and blinks, to indicate where attention is focused on a screen and how it moves between different areas (Mostafa and Gwizdka, 2016; Siegle et al., 2008).
+By annotating areas of interest (AOIs) on results, sentences, or words, eye-tracking allows for detailed analysis of how users allocate attention to attitude-consistent versus inconsistent information, revealing confirmation bias (Nelius et al., 2020).
+Thus, the eye-tracking method was applied to investigate confirmation bias during system interaction in this study (see detailed measurements in Section 4.4)
+Previous research has shown that confirmation bias is widely manifested in seeking behavior, influencing how users interact with information when using traditional search engines.
 Specifically, users may be more likely to select and spend more time reading attitude-consistent results (White, 2013; Suzuki and Yamamoto, 2020), while avoiding information that challenges their prior beliefs (El and Moussaoui, 2022; Knobloch-Westerwick et al., 2015; Dickinson, 2020).
-This behavior may further result in filter bubble and echo chamber (Quattrociocchi, 2017), where like-minded people gather, communicate and interact as an identified community, reinforcing shared beliefs and perspectives (Garrett, 2017).
-To mitigate such biases, HCI and CSCW scholars have explored design strategies in visualization and recommendation systems to encourage user engagement with diverse content.
-One established approach is to increase exposure to opposing viewpoints.
-For instance, Lex et al.
-(2018) mixed tweets and hashtags from various partisan groups, Schweiger et al.
+This behavior may further result in a filter bubble and echo chamber (Quattrociocchi, 2017), where like-minded people gather, communicate, and interact as an identified community, reinforcing shared beliefs and perspectives (Garrett, 2017).
+Advancements in LLM-powered information-seeking have amplified user susceptibility to this bias.
+For example, Sharma et al.
+(2024) compared traditional search systems with conversational search powered by LLMs (with and without references).
+Their results reveal that “participants engaged in more biased information querying” when using conversational search.
+This trend is consistent with prior research in HCI, CSCW, and IR, which shows that users often rely on heuristics to reduce cognitive dissonance — whether due to conflicting beliefs or a mismatch between beliefs and behaviors, when processing complex information (White and Hassan, 2014; Sparrow et al., 2011; White, 2013).
+Azzopardi (2021) argues that this information overload exacerbates related biases, including confirmation bias, anchoring bias, and framing effects.
+Sharma et al.
+(2024)’s study adds to this growing body of evidence, emphasizing thoughtful design considerations for LLM-powered systems, prioritizing guided interactions over simplistic integration.
+2.2.2. Debiasing Strategy in Search Interface Design and System Improvements
+A common Debiasing Strategy is to employ design features in visualization and recommendation systems to help users process diverse content more effectively.
+Though exposing users to diverse perspectives can increase information overload, it remains a crucial first step in mitigating confirmation bias.
+To address this trade-off, researchers have explored ways to make attitude-inconsistent information more visually appealing, better organized, or concise.
+For example, Schweiger et al.
 (2014) used tag clouds to display opposing viewpoints, and Huang et al.
-(2012) presented counterarguments on a separate screen.
-Another approach is to increase the system interactivity and user engagement.
-For instance, Sude et al.
-(2021) added a ”voting” feature for search results, and Rieger et al.
-(2021) applied warning labels like ”this search result may reinforce your opinion,” requiring users to click a button to view them.
-These features help users to be more aware of their information seeking and decision framing process.
+(2012b) presented counterarguments on a separate screen to reduce cognitive interference and encourage more balanced consideration.
+Another approach is to increase the system interactivity or impose cognitive forcing.
+For example, Sude et al.
+(2021) added a “voting” feature for search results, and Rieger et al.
+(2021b) applied warning labels like “this search result may reinforce your opinion,” requiring users to click a button to view them.
+These features help users to be more aware of their information-seeking and decision-framing process.
 Additionally, developers created search systems to balance different search results by learning user existing information activities, such as Schwind et al.
-(2012)’s search system that highlights search result inconsistent with user preferences reflected on their search queries and Jasim et al.
+(2012)’s search system that highlights search results inconsistent with user preferences reflected on their search queries and Jasim et al.
 (2022)’s bias mitigation recommendation model that suggests diverse sentiment review for online products, informed by user past browsing history.
-Aforementioned studies suggest that presenting diverse information in a more vivid way can motivate users to engage with content that differs from their prior views.
-However, it’s important to recognize that these design and system enhancements remain centered on a static, passive model of information consumption, characteristic of traditional search contexts where information is retrieved and displayed rather than dynamically presented.
-While increasing exposure to diverse information through external cognitive forcing functions can increase information diversity, this approach may not inherently enhance user engagement or persuasiveness.
-In contrast, LLM-powered information seeking presents a novel interaction paradigm that enables more diverse, interactive, and dynamic engagement, potentially reducing confirmation bias.
-2.3. Opportunities of LLM Constructive Deliberation on Contrasting Viewpoints
+However, we argue that the aforementioned debiasing strategies remain centered on a static, passive model of information consumption particularly in traditional search experience, where information is retrieved and displayed rather than presented proactively.
+While increasing exposure through system interactivity or cognitive forcing helps foster user understanding of diverse viewpoints, this may not inherently enhance user engagement or persuasiveness.
+In contrast, LLM simulation and role-play (described in Section 2.1) presents a novel interaction paradigm that enables more interactive and dynamic engagement, potentially reducing confirmation bias.
+We introduce this idea along with recent related research in the next section.
+2.3. Constructive Deliberation with LLMs
 Confirmation bias is often associated with polarization effects in which social divides serve to establish and/or reinforce certain viewpoints.
-While exposure to opposing views in a hostile setting can heighten polarization (Bail et al., 2018), constructive deliberation between opposing views can actually promote empathy, tolerance, and sharing ideas, and reduce polarization (Mutz, 2002; Porter and Schumann, 2018; Levendusky and Stecula, 2021; Bächtiger and Parkinson, 2019; Kim et al., 2021).
-Researchers have explored how LLMs can facilitate constructive deliberation and enhance engagement with contrasting viewpoints through mediated discourse.
+While exposure to opposing views in a hostile setting can heighten polarization (Bail et al., 2018), constructive deliberation between opposing views can actually promote empathy, tolerance, and sharing ideas and reduce polarization (Mutz, 2002; Porter and Schumann, 2018; Levendusky and Stecula, 2021; Bächtiger and Parkinson, 2019; Kim et al., 2021).
+Scholars in CSCW developed systems to foster constructive deliberation on opposing viewpoints.
+We found similar design features align closely with those outlined in the previous section.
+By employing similar debiasing strategy mentioned in Section 2.2, the ConsiderIt system, designed by Kriplean et al.
+(2012), “guides people to reflect on trade-offs and the perspectives of others” by presenting pro and con points separately in online debates about U.S.
+state elections.
+Systems that are more relevant to crowd information-seeking, such as Poli (Semaan et al., 2015) and PolicyScape (Kim et al., 2019), allow users to gather information reflecting diverse opinions on social media.
+Additionally, Kim et al.
+(2021) developed StarryThoughts, a system that visualizes diverse online opinions using the metaphor of “starry nights,” aiming to increase user interactivity.
+With the emergence of LLM personas, researchers start applying this to facilitate constructive deliberation.
 From the technical side, Tessler et al.
-(2024) develop the Habermas Machine, a system of LLMs based on Chinchilla which generates statements that express common ground amongst a set of opposing views.
+(2024) developed the Habermas Machine, a system of LLMs based on Chinchilla that generates statements that express common ground amongst a set of opposing views.
 Kambhatla et al.
-(2024) operationalize receptiveness into a computational framework, and use an LLM in combination with the framework to reframe Reddit comments to be more receptive to other users’ comments.
-Other work also studies grounding, or constructing a shared basis of understanding, in promoting positive discussions on opposing views by fine-tuning specific language model (Cho and May, 2020; Shaikh et al., 2024b).
-Researchers in social science and HCI also design LLM-powered tools to facilitate constructive deliberation.
+(2024) operationalize receptiveness into a computational framework and use an LLM in combination with the framework to reframe Reddit comments to be more receptive to other users’ comments.
+Other work investigates grounding a shared basis of understanding in promoting positive discussions on opposing views by fine-tuning specific language model (Cho and May, 2020; Shaikh et al., 2024b).
+Social science and HCI researchers are actively exploring the design of LLM-powered constructive deliberation.
 For example, Argyle et al.
-(2023) aim to improve conversation quality in divisive online political conversations using an AI chat assistant.
+(2023) aims to improve conversation quality in divisive online political conversations using an AI chat assistant.
 Similarly, Zhang et al.
 (2024) designed a multi-agent dialogue system in social media conversations.
-Although no comparison was made with other baseline conditions, they found that participants were influenced either by their own interests or by the system itself to actively explore diverse perspectives, helping to reduce their filter bubbles.
+Although no comparison was made with other baseline conditions, they found that participants were influenced either by their own interests or by the system to explore diverse perspectives, helping reduce their filter bubbles actively.
 Chiang et al.
-(2024) directly use LLM to play devils’ advocate in AI-assisted group decision-making and found that participants were more likely to form an appropriate trust to adopt AI decisions.
-Overall, we see LLMs increasingly being used to reduce polarization and intolerance in ways that might complement LLM-tooling to mitigate related biases.
-Notably, LLM could be designed to simulate personas that take more critical stances according to predefined roles and characters, which might challenge user prior beliefs.
-By presenting a new user experience built around diverse feedback and engagements than traditional search, this approach might help counteract the biased information-seeking behaviors commonly seen in traditional search systems.
-Research Focus of this Study.
-Building on prior research, we aim to thoroughly and objectively assess the effectiveness of LLM-generated personas in reducing confirmation bias within an automated debate context.
-LLM-powered diverse information-seeking via personas demonstrates how LLM agents actively and autonomously can present information and arguments, creating an argumentative experience between humans and AI.
-Unlike traditional information-seeking that is predominantly user-driven, this argumentative experience could foster a more interactive and collaborative exchange of information to reduce both motivational and cognitive confirmation bias.
-While prior studies have examined LLM-generated personas for product recommendations (Park et al., 2023a), reducing filter bubbles through qualitative feedback (Zhang et al., 2024), and generating counterarguments in group decision-making (Chiang et al., 2024), our study objective is to ground the hypotheses (as highlighted above) from a psychological cognitive perspective.
-We designed a multi-persona debate system that presents diverse viewpoints through argumentative discourse on controversial issues.
-By comparing this system with a baseline retrieval-based approach that presents balanced content snippets from opposing perspectives, we evaluate how dynamic, argumentative information presentation influences confirmation bias and belief formation.
-Through eye-tracking analysis capturing users’ involuntary (bottom-up) and goal-directed (top-down) eye movements, we examine how this active, LLM-driven interaction paradigm shapes user engagement.
-In the following sections, we present the system design for the multi-persona debate system (Section 3) and detail the study procedure (Section 4).
+(2024) directly used LLM to play devil’s advocate in AI-assisted group decision-making and found that participants were more likely to form an appropriate trust to adopt AI decisions.
+Research Focus of this Study
+After reviewing prior work, we argue that LLM multi-personas debate creates an argumentative experience to reduce confirmation bias and polarization.
+LLM role-play and simulation offer an engaging user experience that is dynamic and proactive (described in Section 2.1).
+By fostering constructive deliberations (described in Section 2.3), it can mitigate confirmation biases exhibited in traditional and conversational search systems (described in Section 2.2)
+Recent work has explored this emergent user experience, as discussed in Section 2.3.
+However, these studies predominantly validate effectiveness through subjective self-report measures such as verbal accounts and user ratings.
+While such methods yield valuable user insights, they can be complemented and strengthened by incorporating behavioral and cognitive indicators that directly capture confirmation bias, such as eye-tracking data.
+Prior research in human perception and psychology has shown a strong positive correlation between visual attention and confirmation bias (as discussed in Section 2.2).
+Building on this, our work integrates eye-tracking metrics as an objective, real-time measure of attentional patterns, helping to bridge the gap between subjective feedback and quantifiable cognitive behavior.
+Specifically, we developed a multi-persona debate system that leverages LLM to foster critical engagement with controversial topics.
+The system operates by enabling users to create diverse AI personas that debate opposing viewpoints autonomously, encouraging users to synthesize conflicting arguments and mitigate confirmation bias.
+To evaluate this approach, we compare it against a baseline retrieval-based system that passively displays pre-curated, balanced content snippets from opposing perspectives.
+Three corresponding research questions are raised.
+RQ1: What is the overall user experience when conceptualizing controversial topics through a multi-persona debate interaction simulated by LLM?
+RQ2: Compared to the retrieval-based system, how effective is the multi-persona debate system in mitigating confirmation bias?
+RQ3: Compared to the retrieval-based system, how does the multi-persona debate system finally influence user pre-existing beliefs?
+Prior work has predominantly focused on RQ1, with qualitative exploration of RQ2 and tangential implications for RQ3.
+Our study advances this by quantitatively associating confirmation bias reduction with direct belief changes, thereby grounding the efficacy in empirical evidence.
 3. System Design
@@
 Figure 1. Overview of the Multi-Persona Debate System
+[…]
+Overview of the Multi-Persona Debate System
 3.1. Envisioned Use Case
@@
 Figure 2.
-The interface of our multi-persona debate system is shown for the topic “Should human colonize space?” LLM-generated personas are shown at left, which users can freely modify, replace, or delete.
+The interface of our multi-persona debate system is shown for the topic “Should humans colonize space?” LLM personas are shown on the left, which users can freely modify, replace, or delete.
 Once the debate starts, thumbs up/down buttons to the top-right of each argument can be used to bookmark it for the user’s later review, recording whether or not the given argument resonated with the user’s own thinking.
 Bookmarked arguments are collected for the user’s review in the Summary tab.
+[…]
+The interface of the multi-persona debate system
 3.3. Interface Features
-Figure 2 presents the user interface of our LLM debate system. We describe the major interactive features below.
+Figure 2 presents the user interface of the LLM multi-persona debate system. We describe the major interactive features below.
 3.3.1. Controversial Topics Entry or Selection
 At the start, users can either enter a topic, possibly controversial, in the search bar or choose one from the list of suggested topics A.
-The system then begins generating LLM-personas relevant to that topic.
+The system then begins generating LLM personas relevant to that topic.
 3.3.2. Persona Creation
 As displayed in B, each persona is represented by a title and description, along with automatically assigned colors and icons to help users easily identify and differentiate between them (which emerged as a need during early pilot testing).
-We set a predefined number of personas at the beginning (N=5), but users are free to generate or manually create as many personas as they want.
-The specific prompt instructions (described in Section 3.4.1) aim to generate diverse LLM personas that represent a range of viewpoints and social roles on the selected topic.
-This approach aligns with our primary design goal of comprehensive representation.
+We set a predefined number of personas at the beginning (N=5).
+These personas are created automatically by an LLM based on the system prompt: “Given the topic, [TOPIC], create a roundtable debate of different personas to show key perspectives of the topic.” Each persona is associated with its profession and description.
+This prompt instruction (described in Section 3.4.1) generates diverse LLM personas.
+Additionally, to align with our primary design goal of comprehensive representation, users are free to generate or manually create as many personas as they want or edit the existing personas in the interface.
 3.3.3. Interactive Debate Flow
-During the design of this interface, we addressed a critical design consideration regarding the optimal level of inter-persona interaction within each debate round.
+During the design of this interface, we addressed the design consideration regarding the optimal level of inter-persona interaction within each debate round.
 In a structured human debate [^3] https://debatingmatters.com/debate-timing-structure/ where debaters are separated into two teams holding opposing views, responses may either focus on arguments presented by a specific debater or address all previously discussed points made by members from the opposing team.
@@
 3.4.1. Persona Creation
-In creating personas for our multi-persona debate system, prior studies have pointed out several limitations inherent to LLMs (Park et al., 2022, 2023b), such as inconsistent persona generation and formatting.
-While LLMs can generate personas that clearly express their stances on controversial topics, determining how to effectively prompt the LLM to achieve this may not always be straightforward.
-One problem is that assigning specific names to personas, real or fictional, can be problematic because readers may mistakenly believe the persona represents an actual person, as evidenced by prior studies that LLMs would inadvertently reveal real names from their training data (Huang et al., 2022).
-To mitigate misrepresentation and potential ethical concerns, we prompt the LLM to ”generate a title for the persona, not a name”, while enabling users to override this default setting through customization.
-In generating persona descriptions, the LLM may include detailed background information without explicitly clarifying the persona’s stance on the topic, yet explicitly linking persona background to specific stance risks perpetuating stereotypes or creating caricatures of those backgrounds.
+To satisfy our first design objective (i.e., broad representation), we identified two key challenges from prior research and our pilot implementations.
+First, prior work generated personas by constraining them based on different characteristics, such as interest (e.g., hobbies and jobs) and personal traits (e.g., kind, bully) (described in Section 2.1).
+But, we found that this method introduces a tradeoff for our study.
+Overly detailed specifications (e.g., exhaustive lists of traits) risk producing personas that might limit LLMs’ ability to generate nuanced and diverse perspectives.
+Conversely, oversimplified personas with minimum specifications may inadvertently amplify LLMs’ inherent biases, particularly those predisposed to favor dominant cultural or demographic traits (Salminen et al., 2024; Beck et al., 2023).
+This suggests that failing to specify appropriate traits can lead to an imbalanced spectrum of viewpoints, thereby not achieving our goals.
+Another problem is that while assigning specific names, real or fictional, to personas can enhance authenticity and elicit more genuine user responses, it risks misleading readers into perceiving personas as representations of real individuals.
+This concern is amplified by evidence from prior studies showing that LLMs may inadvertently leak real names from their training data (Huang et al., 2022).
+To address diversity limitations, we prompt the LLM to use multiple perspectives to reflect a topic and then generate personas.
+This ensures that even for topics like “Should abortion be banned?” that might favor certain sides, the resulting personas can capture a balanced spectrum of viewpoints, representing both supporting and opposing arguments.
+Additionally, we incorporate existing personas as context when users initiate the generation of a new persona.
+The prompt includes instructions to maintain consistency across persona professions and descriptions.
+To mitigate misrepresentation and potential ethical concerns, we prompt the LLM to “generate a title for the persona, not a name” while enabling users to override this default setting through customization.
+In generating persona descriptions, the LLM may include detailed background information without explicitly clarifying the persona’s stance on the topic, yet explicitly linking persona background to stance risks perpetuating stereotypes or creating caricatures of those backgrounds.
 To address this, we first prompt the LLM to “describe background (instead of their stance)” and then separately prompt it to “start the debate by briefly presenting an argument in support of its stance on the topic.”
-We generate an initial set of personas to cover key perspectives on the issue.
-To enhance perspectival diversity, we incorporate existing personas as context when generating additional ones.
-Our prompt design includes specific instructions to maintain consistency across persona professions and descriptions (see Appendix A for prompt details).
 3.4.2. Debate Generation
-To facilitate the debate, we developed two prompts: one to initiate the debate and another to continue it based on prior debate conversations (see Appendix A for prompts details).
-However, it is important to note that LLMs are constrained by a maximum input length, known as the ‘context window’ (Wang et al., 2024) after which they may start to lose track of earlier portions of the input.
+To facilitate the debate, we developed two prompts: one to initiate the debate and another to continue it based on prior debate conversations.
+However, it is important to note that LLMs are constrained by a maximum input length, known as the ‘context window’ (Wang et al., 2024), after which they may start to lose track of earlier portions of the input.
 This limitation can be problematic in our system, as personas need to reference the entire debate history to provide meaningful responses.
 To overcome this challenge, inspired by Park et al.
-(2023b)’s work, when the debate history exceeds the LLM’s context window, we employ an extractive summarizer that compresses the transcript to fit within the required length while preserving the key points of the discussion.
+(2023b)’s work, when the debate history exceeds the LLM context window, we employ an extractive summarizer that compresses the transcript to fit within the required length while preserving the key points of the discussion.
 Specifically, we use the BERTExtractiveSummarizer (Miller, 2019), which provides a good balance between summarization performance and latency.
@@
 We developed the system UI using Streamlit [^5] https://streamlit.io/, a popular machine learning web development framework that allows development with an emphasis on interactivity and ease of use.
-It also promotes rapid prototyping, which made it easy to iterate on the interface design based on user feedback or emerging requirements.
+It also promotes rapid prototyping, which makes it easy to iterate on the interface design based on user feedback or emerging requirements.
 4. Methods
-To explore how users might engage with the multi-persona debate system and the system effectiveness of reducing confirmation bias, we conducted a controlled, within-subjects study by comparing the multi-persona debate system with a baseline retrieval-based system.
+To explore how users might engage with the multi-persona debate system and the system’s effectiveness in reducing confirmation bias, we conducted a controlled, within-subjects study by comparing the multi-persona debate system with a baseline retrieval-based system.
 In this section, we describe our formal evaluation protocol, including the experimental procedure and tasks (Section 4.1), participant recruitment (Section 4.2), the selection of controversial topics (Section 4.3), and data collection and analysis.
@@
 Figure 3. Flowchart of the experimental procedure.
-As illustrated in Figure  3, the within-subject experiment started with eye calibration and task instructions.
+[…]
+Flowchart of the experimental procedure
+The study followed a within-subjects design to reduce individual variability.
+As each participant interacted with both the baseline system and the multi-persona debate system, this approach allows for direct comparison while controlling for differences between individuals.
+To minimize potential order effects, both the task topics and the sequence of system exposure were randomized across participants.
+As illustrated in Figure  3, the experiment started with eye calibration and task instructions.
 Participants were then randomly assigned to one of two study blocks, Block A or Block B.
@@
 ArgumentSearch is a retrieval-based search engine that identifies articles from the open web and organizes content snippets according to their stance on the user query (detailed description is provided in Appendix C).
+The design of ArgumentSearch explicitly organizes arguments by stance and exposes users to multiple viewpoints from diverse web sources (Daxenberger et al., 2020).
+Compared to traditional search engines, which have proven to inadvertently reinforce confirmation bias (discussed in Section 2.2), the structured presentation of ArgumentSearch makes it a stronger baseline to compare with our multi-persona debate system, enabling a rigorous evaluation of our system effectiveness in mitigating confirmation bias.
 In Block B, participants interacted with our multi-persona debate system as the experimental condition to explore and generate debate content with the given topic.
@@
 Within each block, participants completed four randomized trials assigned with different controversial topics.
-Once they completed the trials in one block, they moved on to the other block and used the alternate system, ensuring they interacted both systems.
+Once they completed the trials in one block, they moved on to the other block and used the alternate system, ensuring they interacted with both systems.
 Each trial consisted of three parts: (i) Pre-task Questions (ii) Exploring the given topic with the assigned interface, and (iii) Post-task Questions.
-In the pre-task questions, we asked participants to explicitly rate their topic familiarity, pre-existing belief (i.e., stance) to the topic, and the confidence for their stance.
-For each trial, participants were presented with a controversial issue to explore the interface.
-After participants completing the task, we again evaluate their current stance to the topic and the confidence score, similar to the pre-task questions.
+In the pre-task questions, we asked participants to explicitly rate their topic familiarity, pre-existing belief (i.e., stance) of the topic, and confidence in their stance.
+For each trial, participants were presented with a controversial topic to explore the interface.
+After participants complete the task, we again evaluate their current stance on the topic and the confidence score, similar to the pre-task questions.
 Upon completion of all eight trials across both systems, participants were required to complete a post-test questionnaire to assess their propensity for confirmation bias.
-Following this, a semi-structured interview was conducted to gain insight into the users’ interactions with the systems, their perception, as well as the usability and feedback.
-Details of the pre-, post-task questionnaires and interview questions are presented in Appendix B.
+Following this, a semi-structured interview was conducted to gain insight into the users’ interactions with the systems, their perception, and the usability and feedback.
+Details of the pre- and post-task questionnaires and interview questions are presented in Appendix B.
 The entire experimental study took approximately 1.5 to 2 hours to complete.
@@
 4.2. Participant Recruitment
-The study was conducted in a usability lab at a university, with $N=40$ participants (19 females, 18 males, and 3 non-binary identity).
-Particpants were recruited using convenience sampling, and were aged between 18 and 34 years (view demographic details in Appendix D).
-They came from a variety of academic disciplines, ranging from applied sciences, natural sciences, engineering, to social sciences, bringing diverse educational backgrounds.
-Participants were pre-screened for native-level English familiarity, 20/20 vision (uncorrected or corrected).
-Upon completion of the study, each participant was compensated.
+The study was conducted in a usability lab at a university, with $N=40$ participants (19 females, 18 males, and 3 non-binary identities).
+Participants were recruited using convenience sampling and were aged between 18 and 34 years (view demographic details in Appendix D).
+They came from a variety of academic disciplines, ranging from applied sciences, natural sciences, and engineering to social sciences, bringing diverse educational backgrounds.
+Participants were pre-screened for native-level English familiarity and 20/20 vision (uncorrected or corrected).
+The study was approved by the Institutional Review Board (IRB), and each participant was compensated with 30 USD upon completion.
 In the data analysis, a total of 156 trials, derived from 39 participants each completing 4 tasks within the multi-persona debate system, were utilized for behavioral (Section 5.1) and eye-tracking analysis (Section 5.2).
@@
 4.3. Selection of Controversial Topics
-These topics were sourced from procon.org [^7] ProCon.org_2024 covered a diverse range of categories, including health, science and technology, economics, and entertainment.
+These topics were sourced from procon.org [^7] ProCon.org_2024 and covered a diverse range of categories, including health, science and technology, economics, and entertainment.
 These topics are formulated as yes/no proposed actions, such as ”Should abortion bans be illegal?” or ”Should space be colonized?”.
@@
 the “con” stance against it.
-We select topic categories to cover a range of important issues with broad societal relevance.
-Each topic was selected to ensure active public discussion, involvement from various groups, reliable information from credible sources, and a balanced view of arguments, without needing much specialized knowledge (view topic details in Appendix 9).
+We chose these topics as they reflect polarized political discourses currently.
+Participants might hold strong prior beliefs about them, making confirmation bias more likely to influence how they process information.
+These topics provide a relevant context for evaluating whether the multi-persona debate system can promote open-minded engagement and effectively mitigate biased attention to attitude-consistent information.
+Each topic was selected to ensure active public discussion, involvement from various groups, reliable information from credible sources, and a balanced view of arguments without needing much specialized knowledge (view topic details in Appendix 9).
 In each block, one topic was selected for each category, resulting in four tasks per block and a total of eight tasks.
+To ensure the quality and stance alignment of the generated arguments, each persona was prompted with structured instructions to reflect its assigned viewpoint on the topic.
+We also conducted post-hoc annotations to verify that each persona consistently expressed its intended stance.
+Personas identified as holding a neutral stance were excluded from the data analysis focused on quantifying patterns of user interaction with attitude-consistent versus attitude-inconsistent arguments.
+In the multi-persona debate system design, users are allowed to freely select personas representing diverse stances.
+To address potential imbalances in stance exposure, we normalized the collected behavioral data, particularly the eye-tracking measures, by accounting for both the number of pro and con personas and the length of the generated debate content.
+This normalization reduces exposure effects and enables more accurate comparisons of user engagement across different viewpoints.
 4.4. Data Collection and Analysis
-Data collection during the experimental procedure incorporated multiple sources: pre-task questions, user interaction logs, post-task questions, system evaluation questionnaires, and interview recordings.
-Eye-tracking data were gathered using the Tobii TX-300 eye-tracker and processed through Tobii Pro Lab  [^8] https://www.tobii.com/products/software/behavior-research-software/tobii-pro-lab, a commercial usability and eye-tracking software that enabled raw gaze data collection, user interaction recording, and preliminary data cleaning for subsequent analysis.”
-Further details about our analysis are provided below.
+Data collection during the experimental procedure incorporated multiple sources: pre-task questions, user interaction logs and eye movements, post-task questions, system evaluation questionnaires, and interview recordings.
+Eye-tracking approach offers objective insights into users’ visual attention and engagement with system content, particularly in assessing how confirmation bias manifests during interaction.
+Eye-tracking data were gathered using the Tobii TX-300 eye-tracker and processed through Tobii Pro Lab  [^8] https://www.tobii.com/products/software/behavior-research-software/tobii-pro-lab, a commercial usability and eye-tracking software that enabled raw gaze data collection, user interaction recording, and preliminary data cleaning for subsequent analysis.
+Eye-tracking data and behavior logs were filtered to exclude segments involving task scenario reading and pre/post-task questionnaires, retaining only the periods of active interface interaction.
+For analysis, we removed low-quality and non-reading eye-tracking data by discarding fixations shorter than 100 ms or longer than 1500 ms.
+After preprocessing, statistical assumptions were assessed for all analyses reported in the paper.
+Additionally, normality tests (mainly Shapiro-Wilk and Q-Q plots) were performed on the dependent variables, including eye-tracking measures and user belief change data, followed by the use of parametric or non-parametric tests.
+Further analysis details are provided below.
 4.4.1. RQ1: What is the overall user experience when conceptualizing controversial topics through a multi-persona debate interaction simulated by LLM?
-To categorize user interaction patterns with our LLM debate system, we conducted a clustering analysis using the following behavioral data:
+To answer this question, we employed both quantitative and qualitative data analysis.
+We first analyzed the interaction patterns that participants developed when using the system and correlated them with qualitative data.
+A clustering analysis was used to uncover interaction patterns based on the following behavioral data :
 Debate rounds:
@@
 Since we compare user experience between two different interfaces, our qualitative themes focus on factors influencing participant comparisons between the A/B systems used.
-We present these findings in Section 5.1.
+Two rounds of qualitative coding were conducted, each performed by different researchers.
+In the first round, two researchers employed bottom-up thematic open coding without referencing the quantitative data, ensuring an unbiased data exploration.
+In the second round, a third researcher performed thematic axial coding based on the clustered groups.
+The above analysis resulted in two main findings to answer the first question: (1) an integrated result combining quantitative and qualitative insights that illustrate participants’ behavioral patterns, and (2) a set of qualitative themes that capture overarching user behaviors when engaging with the tool.
+We present them in Section 5.1.
 4.4.2. RQ2: Compared to the retrieval-based system, how effective is the multi-persona debate system in mitigating confirmation bias?
-A primary research objective is to evaluate the system’s effectiveness in mitigating confirmation bias in information behavior, particularly in reducing selective exposure to attitude-consistent information.
-Through eye-tracking data analysis, we examine engagement patterns with content that diverges from users’ prior beliefs.
-Drawing on confirmation bias literature (described in Section 2.2), we hypothesize that increased temporal engagement with reading opposing viewpoints may indicate reduced confirmation bias through enhanced exposure to differing perspectives.
-Additionally, we incorporate two key factors in evaluating the above hypothesis based on our study design.
-First, we considered the impact of users’ general tendency toward confirmation bias as a psychological disposition (described in Section 2.2), to understand how this human factor may affect the system’s effectiveness.
-The second factor is user topic familiarity, since we intentionally selected different social topics (described in Section 4.3).
-We thus formulate two further sub-questions regarding how users’ eye movements might differ based on these factors.
+A primary research objective is to evaluate the system effectiveness in mitigating confirmation bias, particularly in reducing selective exposure to attitude-consistent information.
+Drawing on cognitive literature studying confirmation bias user engagement — specifically, heightened attention to content aligning with users’ prior beliefs — serves as a cognitive manifestation of confirmation bias (Betsch et al., 2001; Jonas et al., 2008) (the Motivational aspect of confirmation bias, described in Section 2.2).
+To investigate this, we leverage eye-tracking data to analyze engagement patterns with attitude-inconsistent information (i.e., viewpoints diverging from users’ existing stances).
+We hypothesize that prolonged dwell times on such content may signal reduced confirmation bias, as extended exposure to opposing viewpoints encourages cognitive reevaluation.
+Additionally, we incorporate two key independent variables (IV) in evaluating the above hypothesis based on our study design.
+First, we considered the impact of users’ general tendency toward confirmation bias as a psychological disposition (the Cognitive aspect of confirmation bias, described in Section 2.2), investigating how this human factor affects the system effectiveness.
+The second variable is user topic familiarity since we intentionally selected different social topics (described in Section 4.3).
+Corresponding eye-tracking data is our dependent variable (DV). We thus formulate two further sub-questions regarding how users’ visual attention might differ based on these factors.
 RQ2.1:
@@
 RQ2.2:
-How does use of our LLM debate system vs.
+How does the use of LLM multi-persona debate system vs.
 the baseline system impact user engagement with content?
@@
 These measures are defined as follows:
-Topic familiarity:
+IV: Topic familiarity:
 Users are categorized into five levels of familiarity based on a 5-point Likert scale, as collected in the pre-survey.
-Confirmation biases tendency:
+IV: Confirmation biases tendency:
 Participant levels of confirmation bias were measured using Berthet (2021)’s questionnaire.
@@
 To categorize participant confirmation bias tendencies, we divided the scores into three equal groups, classifying them into low, medium, and high.
-Eye-tracking:
-Two metrics, including total fixation duration and fixation counts, were utilized to understand user engagement.
+DV: Eye-tracking:
+Two metrics were utilized to understand user engagement, including total fixation duration and fixation counts.
 Fixations were filtered to include only good-quality fixations, lasting between 100ms and 2000ms.
 Additionally, both fixation duration and counts were normalized relative to the length of the content.
-Areas of interests:
-We also label AOIs to identify participant selective exposure to attitude-consistent versus attitude-inconsistent information.
+DV: Areas of interests:
+We label AOIs to identify participant selective exposure to attitude-consistent versus attitude-inconsistent information.
 For example, if the content supports the given topic and the participant also holds a supportive view, then the content would be labeled as attitude-consistent information.
-Fixations falling into these AOIs represent how user engage with the attitude-consistent information.
+Fixations falling into these AOIs represent how users engage with the attitude-consistent information.
 Table 1 presents how we annotate attitude-consistent and attitude-inconsistent AOIs based on user pre-task belief and the stance of the content.
@@
 Table 1. Belief consistency of areas of interest (AOIs) in relation to pre-task belief and content stance.
-We present the results for RQ2 in Section 5.2, focusing on user engagement of the content.
+We present the results for RQ2 in Section 5.2
 4.4.3. RQ3: Compared to the retrieval-based system, how does the multi-persona debate system finally influence user pre-existing beliefs?
-We adopt a similar approach as with RQ2. Correspondingly, we formulate two sub-research questions relevant to topic familiarity and user confirmation bias tendency:
+We adopt a similar approach as with RQ2.
+Note that a reduced confirmation bias, as manifested in user engagement of attitude-inconsistent information, might not finally result in user belief changes.
+To validate this relationship, we formulate two sub-research questions relevant to the prior IV: topic familiarity and user confirmation bias tendency:
 RQ3.1:
@@
 RQ3.2:
-How does our LLM debate system, compared to the baseline system, impact user final belief changes?
+How does the LLM multi-persona debate, compared to the baseline system, impact user final belief changes?
 Additionally, does this effect differ based on user levels of topic familiarity and their confirmation bias tendency?
+To address both questions, we measure user belief changes (i.e., the difference between user pre- and post-belief) as another DV.
 User pre- and post-belief were collected separately in the pre- and post-task questions.
-We calculated three types of belief changes, including non-directional, bi-directional, and weighted belief changes based on user confidence, defined as follows:
-Belief change:
+We calculated three types of belief changes, including non-directional, directional, and weighted belief changes based on user confidence, defined as follows:
+DV: Belief change:
 Users are categorized into two groups of belief changes — changed vs.
-unchanged — based on the difference of belief measurement for each topic.
+unchanged (Equation 1) — based on the difference of belief measurement for each topic.
 User belief is measured based on a 5-point Likert scale (i.e., very supportive, somehow supportive, neutral, somehow opposed, and very opposed), collected in the pre- and post-task questions.
-Bi-directional belief change:
-Users are categorized into three groups of belief changes — less supportive, more supportive, and unchanged — showing the directional belief change for each topic, as calculated based on the previous 5-point Likert scale.
-Weighted belief change:
-We calculate the stance shift weighted by the average confidence level before and after using each system (Equation 1) [^9] In prior behavioral studies by Norman and Price (2015), relative confidence of people’s subjective responses were evaluated using average measures.
+DV: Directional belief change:
+Since the prior belief change does not indicate whether users become more supportive or opposed to the topic, we use a directional measure to differentiate both.
+We categorize the directional belief change into — less supportive, more supportive, and unchanged (Equation 2).
+DV: Weighted belief change:
+Furthermore, we calculate the belief change weighted by the average confidence level before and after using each system (Equation 3) [^9] In prior behavioral studies by Norman and Price (2015), relative confidence of people’s subjective responses were evaluated using average measures.
 Building on this, we evaluate participants’ relative confidence in their beliefs about the topics by calculating the average confidence level before and after the intervention..
-This score reflects both the direction and strength of the belief change.
+This score reflects both the direction and strength of changes.
 User belief confidence is also measured by a 5-point Likert scale in the pre- and post-task questions.
-(1) $\frac{(\textnormal{Belief}_{\textnormal{after}}-\textnormal{Belief}_{\textnormal{before}})(\textnormal{Confidence}_{\textnormal{after}}+\textnormal{Confidence}_{\textnormal{before}})}{2}$
+(1) $Belief\ change=\text{sgn}(|Belief_{after}-Belief_{before}|)$
+(2) $Belief\ change_{directional}=\text{sgn}(Belief_{after}-Belief_{before})$
+(3) $Belief\ change_{weighted}=\frac{\text{(}Belief_{after}-Belief_{before})(Confidence_{after}+Confidence_{before})}{2}$
 Results for RQ3 are detailed in Section 5.3, which discusses belief changes.
@@
 5.1. RQ1: User Experience in Conceptualizing Controversial Topics through Multi-Persona Debate
+In this section, we present two categories of findings: (1) a mixed-methods result combining quantitative and qualitative data to reveal participants’ behavioral patterns (Section 5.1.1) and (2) qualitative themes exploring participants’ motivations for editing personas (Section 5.1.2) and their perceptions of the generation quality (Section 5.1.3).
+For brevity, we refer to the LLM-based multi-persona debate system simply as “the debate system” in the following results.
+Similarly, the baseline retrieval-based system is simplified as “the baseline system”.
+5.1.1. Clustered behavioral patterns
 From the behavioral data collected in the debate system (as described in Section  4.4), users generated an average of 1.865 debate rounds per topic.
@@
 Users made minimal edits (0.187 personas), clicked thumbs up/down 2.628 times to bookmark content, and accessed the bookmark tab only 0.121 times on average.
+Participants spent an average of 4.82 minutes interacting with and reading the debate content.
 Full descriptive data are provided in Table  6.
@@
 We characterize a usage pattern for each group as follows:
-5.1.1. Active observers (Group 0)
-Participants in this group frequently used thumbs up / down features to express their opinions while observing the debate, and often reviewed the bookmarked content.
+Active observers (Group 0) Participants in this group frequently used thumbs up / down features to express their opinions while observing the debate, and often reviewed the bookmarked content.
 They typically engaged in more than two debate rounds and showed moderate involvement in selecting and editing personas, balancing supportive and opposing viewpoints.
@@
 It is marked by frequent use of feedback features, revisiting bookmarked content, and a balanced consideration of both supportive and opposing perspectives.
-5.1.2. Multi-round explorers (Group 1)
-Participants in this group were inclined to generate more rounds of debate (up to five) but spend less time reading and digesting arguments presented.
+Multi-round explorers (Group 1) Participants in this group were inclined to generate more rounds of debate (up to five) but spend less time reading and digesting arguments presented.
 Notably, we observed a preference for selecting personas that agreed with the given topic.
@@
 This pattern reflects a strategy of building understanding through repetition and confirmation, with a gradual move toward considering alternative viewpoints as participant knowledge grows throughout the debate.
-5.1.3. Critical readers (Group 2)
-Participants in this group exhibited the highest level of engagement with the system.
+Critical readers (Group 2) Participants in this group exhibited the highest level of engagement with the system.
 Compared to other two groups, they experienced the longest debate durations, selected more new personas, and frequently edited these personas.
@@
 We present these findings, along with further analysis in the Appendix E.
+5.1.2. Motivations for persona edits
+From the post-task interviews, we found three additional factors that participants frequently mentioned in selecting and revising personas, including persona expertise, individuality, and the overall diversity of personas when selecting and editing them to construct the debate.
+These factors were closely tied to how participants conceptualized different controversial topics being discussed.
+First, many participants preferred selecting both expert and individual personas.
+Expert personas were more trusted to provide factual information, evidence, and other logical and reliable arguments (P11, 14-15, 22, 24, 33, 36, 38-39).
+P39 explained that for the topic of “TikTok ban,” experts could provide opinions rather than emotional responses “Parents who are concerned about their kid is great, but I don’t expect them to have all of the data necessary to make an informed decision.
+I thought there would be experts discussing things logically and providing evidence-based information, as opposed to information driven by emotions.” In contrast, individual personas could provide direct and authentic arguments (P11, 22, 24-25, 33, 38).
+For example, P38 said, “Whether TikTok should be banned from school.
+I really want people who are in the school specifically because it’s important to see the people who are actually being affected to know their pain points.
+So that’s why I picked the teacher, the student, and the administrator.
+Additionally, as participants were aware that potential bias could underlie different personas, some aimed to create a balanced representation of opinions by selecting personas that supported different sides of the topic.
+On the one hand, this approach could be an effective strategy for generating counterarguments (P14, 24, 26, 28, 33, 35-36), as P36 explained, “I tried to balance the number of people who supported the topic with those who opposed it, to sort of weigh which perspective truly aligns with my own thinking and against it.” On the other hand, some would deliberately balance opinions that might be disproportionately represented on one side in real-world discussions (P22, 24, 33).
+P33 pointed out that “If I’ve been hearing a lot of one point of view, I’ll try to keep generating a persona until I find one that opposes it.
+5.1.3. Perceptions of generation quality
+Participants employed two main strategies to evaluate the quality of the generated content.
+First, they assessed whether the persona arguments, opinions, and use cases aligned logically with their customized prompts (e.g., political alignment, expertise).
+Second, they examined the concreteness, reasonableness, and creativity of the content.
+Many of them emphasized that this dual-level evaluation helped them further assess the reliability of the multi-persona debate system, particularly given their awareness of potential LLM hallucinations.
+For example, participants prioritized arguments supported by contextual background knowledge.
+When content lacked logical coherence, they often dismissed it as unconvincing and retained their original stance (P19, 29, 30).
+P30 explained that “I was really curious to see really compelling arguments.
+When I saw they mostly centered on a trickle-down economy [Topic: ‘Corporate income tax raise’], which I think there’s a lot of evidence against, I didn’t find them strong enough to change my opinion on the topic.
+Repetitive or irrelevant debate content further diminished perceived usefulness.
+For example, some participants criticized exchanges that failed to address the core topic across multiple rounds (P13, 17, 21, 26, 35-36, 39-40).
+This issue extended to the baseline retrieval system, where irrelevant sourced materials, even when citations were provided, make them feel less useful (P13, 22, 24, 27-28, 39).
+Conversely, participants perceived the system as highly useful when debates met three conditions: (1) the content fostered deeper engagement with the topic (P14, 28); (2) it introduced novel, perspective-shifting knowledge (P11–12, 14–15, 28, 30); or (3) it provoked critical self-reflection through creative or unexpected arguments (P14, 24, 34).
+For example, P30 revised the stance on animal testing after encountering arguments highlighting non-animal technological alternatives, noting: “The argument that we have technology capable of conducting similar tests and obtaining comparable data without involving or harming animals was something I hadn’t considered initially, and it shifted my stance.” Similarly, P24 reconsidered their opposition to violent video games when debates emphasized their potential benefits, “we aren’t inherently violent, and the inability to express aggression doesn’t mean we shouldn’t play.
 5.2. RQ2: Comparison of User Engagement between Retrieval-based and Multi-persona Debate System
@@
 Table 3.
-Comparison of Interface Measurements for Attitude-consistent and Attitude-consistent Conditions.
+Comparison of total fixation counts and fixation durations in the interfaces for Attitude-consistent and Attitude-consistent Conditions.
 The significance levels are represented as p-value, 0 ¡ ‘***’ ¡ 0.001 ¡ ‘**’ ¡ 0.01 ¡ ‘*’ ¡ 0.05.
@@
 Comparison of total fixation duration on attitude-consistent and inconsistent information, displayed in box plots, between baseline and debate systems, analyzed across varying levels of topic familiarity and confirmation bias tendency levels.
+[…]
+Comparison of total fixation duration on attitude-consistent and inconsistent information
 System comparison.
@@
 […]
-The ANOVA analysis conducted above on content-belief consistency but separately with topic familiarity and confirmation bias tendency reveal a consistent result: there is a clear difference on whether participants reading content consistent or inconsistent with their attitude when using the debate system, as reflected in their eye movements, but no difference found in the baseline.
-While topic familiarity yields no difference in both systems, confirmation bias tendency resulted in significant difference on the baseline but not the debate system.
-In combined with the system comparison results shown in the paired t-test (Table 3), we can conclude that the debate system effectively reduces selective exposure and increases user engagement with attitude-inconsistent content, while the baseline system activates confirmation bias tendencies, as evidenced by distinct eye movement patterns.
 5.2.2. Effects of bias mitigation across user behavior patterns
-Figure 5. Comparison of normalized fixation duration and fixation counts on attitude-consistent and inconsistent information across user behavior groups.
+Figure 5. Comparison of normalized fixation duration and fixation counts on attitude-consistent and inconsistent information across user behavior groups while using the debate system.
 […]
-These results suggest that there is significant difference in the eye movements between behavior clusters within the debate system, As content-belief consistency did not significantly influence fixation counts, this suggests that user engagement levels were more closely tied to their behavior patterns when using the debate system rather than whether participants would read content associated with their prior beliefs.
-Additionally, as post-hoc analysis revealed that Critical Readers (group 2) had significantly higher fixation counts compared to the other groups, this suggests that Critical Readers were more visually engaged with the content.
-By associating with qualitative results, describing how Critical Readers actively engaged with various personas and generated content (Section  5.1), we can conclude that certain behavioral pattern, i.e., Critical Readers, show a clear sign of engagement demonstrated in their eye movement data.
+Comparison of normalized fixation duration and fixation counts on attitude-consistent and inconsistent information across user behavior groups while using the debate system
+[…]
 5.3. RQ3: Comparison of User Belief Changes between Retrieval-based and Multi-persona Debate System
 In this section, we first examine whether participant belief changed or remained unchanged across two systems and the influence of topic familiarity and confirmation bias tendency (Section 5.3.1, similar to our approach for RQ2).
-Afterward, we explore in greater depth in terms of bidirectional belief changes, i.e., whether participants became more or less supportive, because this helps suggest whether there is a shift toward increased or decreased polarization (Section 5.3.2).
-As the belief change data does not follow a normal distribution, we applies non-parametric test for analysis.
+Afterward, we explore in greater depth in terms of directional belief changes, i.e., whether participants became more or less supportive (Section 5.3.2).
+As the belief change data does not follow a normal distribution, we apply a non-parametric test for analysis.
 5.3.1. Effects of non-directional belief changes across systems and human factors
@@
 However, there is a less obvious trend shown in confirmation bias tendency.
-We conducted a generalized linear mixed-effects model analysis to quantify the statistical effects (model details were presented in Appendix F.2).
+We conducted a generalized linear mixed-effects model analysis to quantify the statistical effects (model details were presented in Appendix E.3).
 The overall result demonstrates that there is a statistically significant difference across the various levels of topic familiarity and confirmation bias tendency on whether participants changed belief or remained unchanged, while the difference is not shown in interface conditions.
-We briefly describe the mixed effects for each variable in details as follows.
+We briefly describe the mixed effects for each variable in detail as follows.
 Figure 6.
@@
 However, no such trend was evident across different levels of confirmation bias.
+[…]
+Percentage stacked belief changes bar charts between baseline vs. experiment across topic familiarity and confirmation biases tendency
 System comparison: The system variable has an estimated positive coefficient of $0.2474$ ($SE=0.2565$, $z=0.964$ , $p=0.3349$).
@@
 Confirmation bias tendency Confirmation bias tendency shows a statistically significant positive effect on belief change with an estimated coefficient of $0.4338$ ($SE=0.1873$, $z=2.315$ , $p<0.05$).
-This means that participants with higher levels of confirmation bias tendency are more likely change their beliefs, which is counterintuitive to our understanding.
-By examining the different proportions of belief changes across CB groups, we found that the low tendency group had the highest individual percentage of belief change (25.6%), but the combined belief change percentages of the medium and high tendency groups (19.1% + 20.6% = 39.7%) were greater than that.
-This cumulative effect may indicate that participant belief changes might not operate neatly into a three-level continuum of confirmation bias tendency.
-Correlation among three variables: As parts of the generalized fixed-effect results, there is no moderate correlation among group interface, topic familiarity, and confirmation bias tendency, reflecting that these variables operated relatively independently in influencing belief change.
+This means that participants with higher levels of confirmation bias tendency are more likely to change their beliefs, which is counterintuitive to our understanding.
+In a closer examination of belief change patterns across confirmation bias groups, we found a non-linear relationship.
+For example, the low tendency group demonstrated the highest individual percentage of belief change (25.6%), followed by a decrease in the medium tendency group (19.1%) and then a slight increase in the high tendency group (20.6%).
+Although the low tendency group showed the highest individual percentage, the combined effect across medium and high groups (39.7% total) exceeded that of the low group, which explains the positive coefficient in our linear model.
+Correlation among three independent variables: As parts of the generalized fixed-effect results, there is no moderate correlation among group interface, topic familiarity, and confirmation bias tendency, reflecting that these variables operated relatively independently in influencing belief change.
 5.3.2. Effects of directional belief changes across systems and human factors
-We examined how system conditions influence participant bi-directional belief changes, weighted by their confidence level, and explored this variance on topic familiarity and confirmation bias tendency.
+We examined how system conditions influence participant directional belief changes, weighted by their confidence level, and explored this variance on topic familiarity and confirmation bias tendency.
 Measurements Retrieval-based Multi-persona debate Wilcoxon
@@
 Table 4.
-Bi-directional belief change between baseline and experiment condition: a larger number of participants became less supportive after using the experiment system compared to the baseline.
-The weighted change for the experiment was significantly lower than that of the baseline, suggesting that the experiment is more likely to lead participants to become less supportive of their initial stance.
+Directional belief change between baseline and experiment condition: a larger number of participants became less supportive after using the experiment system compared to the baseline.
+The weighted change for the experiment was significantly lower than that of the baseline, suggesting that the experiment is more likely to lead participants to become less supportive of the stance.
 The significance levels are represented as p-value, 0 ¡ ‘***’ ¡ 0.001 ¡ ‘**’ ¡ 0.01 ¡ ‘*’ ¡ 0.05
 Figure 7.
-Bidirectional belief changes box plots for the baseline and experimental interfaces across topic familiarity and confirmation biases tendency: Over both factors, while many participants did not alter their beliefs, which reduced the overall magnitude of changes in the box plots, for participants who did change their beliefs, the multi-persona debate system was more likely to lead participants to become less supportive than the baseline.
-System comparison: By mapping bi-directional belief changes — less supportive, more supportive, and unchanged — on both systems (Table 4), we found that although many participants did not change their beliefs ($Retrieval_{unchanged}$ = $68.2\%$ , $Debate_{unchanged}$ = $62.5\%$), more participants became less supportive after using the debate than the baseline ($Debate_{less}$ = $24.4\%$ $Retrieval_{less}$ = $11.8\%$).
+Directional belief changes box plots for the baseline and experimental interfaces across topic familiarity and confirmation biases tendency: Over both factors, while many participants did not alter their beliefs, which reduced the overall magnitude of changes in the box plots, for participants who did change their beliefs, the multi-persona debate system was more likely to lead participants to become less supportive than the baseline.
+[…]
+Directional belief changes box plots for the baseline and experimental interfaces across topic familiarity and confirmation biases tendency
+System comparison: By mapping directional belief changes — less supportive, more supportive, and unchanged — on both systems (Table 4), we found that although many participants did not change their beliefs ($Retrieval_{unchanged}$ = $68.2\%$ , $Debate_{unchanged}$ = $62.5\%$), more participants became less supportive after using the debate than the baseline ($Debate_{less}$ = $24.4\%$ $Retrieval_{less}$ = $11.8\%$).
 Additionally, by conducting Wilcoxon sign-ranked test on the weighted belief change (in Table 4, the last row, where the mean value ¡0 indicates becoming less supportive and ¿0 indicates becoming more supportive), the weighted belief change for the debate was significantly lower than that of the baseline ($Debate_{weightedchange}$ = $-0.34$ $Retrieval_{weightedchange}$ = $0.30$, $X^{2}=1040.5$, $p<0.05$).
@@
 However, among participants who did change their beliefs, there is a lowered central tendency (median and the interquartile range) in the experiment compared to the baseline.
-This expands on earlier findings of significant difference across levels of topic familiarity (Section 5.3.1), demonstrating that, for participants who finally changed their belief, the debate system led them to become less supportive across different familiarity levels.
-Additionally, we also found the weighted belief change vary depending on different topics (additional analysis and findings are presented in Appendix F.3).
+This expands on earlier findings of significant differences across levels of topic familiarity (Section 5.3.1), demonstrating that, for participants who finally changed their belief, the debate system led them to become less supportive across different familiarity levels.
+Additionally, we also found the weighted belief change varies depending on different topics (additional analysis and findings are presented in Appendix E.4).
 Confirmation bias tendency: By visually presenting participants’ bidirectional belief changes across different levels of confirmation bias tendency, we found a similar trend observed in the topic familiarity (Figure 7, the two plots on the right).
 While many participants did not change their beliefs, reducing the overall magnitude changes as presented in the box plot, for participants who did changed their beliefs, the debate system made user less supportive across three levels of confirmation bias tendency, whereas the baseline tended to increase it.
-By combining the statistical results in both Section 5.3.1 and 5.3.2, we learned that while the system conditions did not yield a statistically significant difference on participant non-directional belief changes (i.e., whether participant belief changed or remained unchanged), for directional belief changes, there was a clear distinction.
-The debate system was more likely to cause participants to become less supportive of their initial stance, whereas the baseline tended to reinforce prior beliefs, making them more supportive.
-Additionally, user belief changes exhibited a statistically significant increase as topic familiarity decreased.
-While a statistically significant difference was also observed with confirmation bias tendencies, the trend did not fully align with our hypothesis.
-This discrepancy may be due to the three-level classification of bias tendency scores, indicating a need for further investigation in future studies.
 6. Discussion
-6.1. Enhancing Interaction
-While our LLM debate system supports several forms of user interaction (e.g., topic selection, persona revision or creation, argument bookmarking, and controlling debate duration), participants offered several ideas to further increase interaction and engagement.
-Specifically, they sought greater ability to actively participate in and influence the debate.
-For example, some participants suggested use of the search bar to allow them to inject their own opinions into an ongoing debate (P2-3).
-Similarly, some participants misunderstood the thumbs up and down bookmarking icons as allowing them to influence the ongoing debate and expressed a desire for this feature when realizing it was not supported (P26).
-Others requested the ability to highlight specific quotes from a persona content and prompt the persona to provide further information or supporting evidence (P25-26, 28, 39).
-A fact-checking feature was suggested to show the most recent evidence from the web for user-highlighted content, aiding in verification of arguments (P39).
+6.1. Summary of Key Findings
+In this section, we synthesize key findings from Section 5 that address each research question, drawing on both quantitative evidence and qualitative insights.
+We also contextualize these findings within the existing literature, examining how they align with or diverge from previous work on technology-facilitated critical thinking (as described in Section 2.3) and approaches to mitigating confirmation bias (Section 2.2).
+6.1.1. Quantitative insights
+Regarding participant experience with the debate system (RQ1), we identified three distinct behavioral patterns through cluster analysis of user interactions (Section 5.1.1).
+Our approach differs from Zhang et al.
+(2024), who also implemented LLM multi-persona debates but categorized user behaviors based solely on whether users modified personas or accepted system defaults.
+Their analysis results in three user groups: interest-driven, system-driven, or mixed.
+Our classification incorporates a more comprehensive set of interaction metrics and more importantly correlates these behavioral clusters with participant visual attention (Section 5.2.2).
+By capturing differences in how users engage with diverse perspectives, personalize personas, and manage their reading attention, this approach supports a more nuanced understanding of user behavior using multi-persona debate and its role in reducing confirmation bias.
+Regarding which systems are more effective in reducing user confirmation bias (RQ2), our statistical analysis (described in Section 5.2.1) reveals that participants paid more attention to attitude-inconsistent content when using the debate system compared to the baseline system.
+Regarding the two additional independent variables — topic familiarity and confirmation bias tendency — we found that while topic familiarity did not affect participants’ visual attention on either system, confirmation bias tendency significantly influenced visual attention in the baseline system.
+These findings suggest that the debate system effectively reduces the Motivational aspect of confirmation bias — increased engagement with attitude-inconsistent content, whereas the baseline system activates the Cognitive aspect of confirmation bias — confirmation bias tendencies.
+In terms of the system influence on users’ final belief changes (RQ3), our statistical analysis (Section 5.3) indicates that system conditions did not produce a statistically significant effect on whether participants changed their beliefs (non-directional change).
+However, there was a notable difference in the direction of changes — more or less supportive to the initial stance on the topic.
+For example, the debate system was more likely to decrease support, while the retrieval system tended to reinforce supportive beliefs.
+Further analysis revealed that these changes varied by topic (see Appendix E.4).
+The retrieval system increased participant support for ‘Animal Testing’ and ‘Space Colonization,’ whereas the debate system decreased support for ‘Federal Corporate Income Tax Raise’ and ‘Violent Video Game Effect on Youth Violence.’ Additionally, there was an overall trend of increased support for ‘Universal Basic Income.’ Moreover, we found that if users were more familiar with the topic, it is less likely for them to change belief.
+Reflecting user qualitative responses (Section 5.1.2 and 5.1.3), we assume that the effectiveness of each system in shifting beliefs depends heavily on the nature of the topic and user pre-conception.
+Compared to prior work, few studies have examined the role of topic differences and explicitly measured user familiarity when evaluating system effectiveness of fostering critical thinking or reducing confirmation bias.
+Most recent work also relies on one or two topics (Kim et al., 2021; Kriplean et al., 2012; Faridani et al., 2010).
+By incorporating multiple topics and accounting for user familiarity, our study offers a more nuanced understanding of how passive information presentation (i.e., retrieved content snippets) versus active argumentative formats (i.e., multi-persona debate arguments) interact with users’ pre-existing beliefs.
+Our study not only highlights the variability in belief shifts across topics but also underscores the need for future research to account for topic-specific effects when designing and evaluating similar systems.
+6.1.2. Qualitative insights
+Participant qualitative responses complement the behavioral patterns found in the quantitative data, further demonstrating that multi-persona debate can effectively serve as technological interventions to promote critical thinking.
+Compared to prior work, we identify similar qualitative factors but with nuanced differences.
+As highlighted in previous studies, identity plays a crucial role in determining user trust toward diverse opinions.
+For example, research utilizing crowdsourced opinions found that participants were concerned about the factuality, authority, and reliability of information sources (Kriplean et al., 2012; Kim et al., 2019).
+Similar concerns arise in our participant feedback regarding the expertise and individuality of LLM personas (Section 5.1.2), as well as the logical robustness and factuality of generated debates (Section 5.1.3).
+Additionally, prior work suggests that identity could create a trust paradox (Kim et al., 2021).
+For example, when users encounter opinions, they often rely on demographic cues to assess an author’s relevant experience and personal stake, which can enhance the perceived credibility of the information.
+However, this evaluative process also risks triggering representativeness heuristic, where users may inaccurately associate certain demographic characteristics with stereotypical viewpoints.
+LLM multi-persona debates carry a risk of reinforcing stereotypes, but this can be partially addressed through carefully crafted persona descriptions that offer depth and context without referencing real individuals.
+The debate system enables participants to customize persona traits through iterative prompting.
+This editing process not only allows users to correct potential misalignment between their real-world perspectives and the arguments generated by the LLM, but also encourages a more reflective evaluation by slowing down the judgment process — helping to counteract the influence of the representativeness heuristic.
+While this cannot fully eliminate inaccurate associations between stereotypical viewpoints and demographic characteristics, it helps mitigate the compounding effects of the representativeness heuristic and confirmation bias, which often reinforce each other in decision-making (Newkirk, 2014).
+In summary, the multi-persona debate system significantly reduces confirmation bias by encouraging users to engage more deeply with content that challenges their initial beliefs, as supported by eye-tracking data and behavioral measures.
+In the following sections, we present additional findings and reflections that, while beyond the primary scope of this study, are closely aligned with directions for future research.
 6.2. Balancing Divergent Viewpoints and Convergent Solutions
-While participants appreciated the system for broadening their understanding of controversial topics through multiple viewpoints, many expressed a desire for more concrete solutions to address the social issues being debated.
-For example, some participants preferred concrete answers over the exchange of opinions.
-P13 noted personas“tend to stop at potential angles for the topic, but they don’t dive much deeper into more specific details.” P29 said “the chat didn’t address specific concerns and I couldn’t find a clear solution.” This led us to further reflect on when a system should offer diverse perspectives vs.
+While participants appreciated the system for broadening their understanding of controversial topics through multiple viewpoints, many expressed a desire for concrete solutions to address the social issues being debated.
+Thus, in this section, we reflect on participants’ qualitative feedback and explore how to extend the technology-facilitated critical thinking tool beyond viewpoint exploration toward a solution-oriented space.
+During our interview, some participants preferred concrete answers over the exchange of opinions.
+P13 noted personas“tend to stop at potential angles for the topic, but they don’t dive much deeper into more specific details.” P29 said “the chat didn’t address specific concerns and I couldn’t find a clear solution.” This preference for concrete solutions may stem from participants’ increased exposure to opposing viewpoints, which could heighten cognitive dissonance.
+While no participants explicitly mentioned experiencing cognitive dissonance, such as complaints about divergent viewpoints, their repeated inquiries about clear resolutions might suggest an underlying discomfort with unresolved tensions in the discussion.
+These observations led us to further reflect on when a system should offer diverse perspectives vs.
 more concrete answers.
-Providing convergent viewpoints can help spur decision-making but may also lead to automation bias in users too readily accepting system answers, something we sought specifically to avoid in our design.
+Note that providing convergent viewpoints can help spur decision-making but may also lead to automation bias in users too readily accepting system answers, something we sought specifically to avoid in our design.
 Bearing a more critical and constructive stance, we consider there are different design opportunities to balance divergent viewpoints with convergent solutions effectively.
@@
 Such bias often stems from the limitations of training data that incorporate existing social inequalities, stereotypes, or incomplete views of the population (Christian, 2020).
-Recent work has examined these misalignment across misinformation (Neumann et al., 2024), climate change (Santurkar et al., 2023), and politics (Taubenfeld et al., 2024).
+Recent work has examined these misalignments across misinformation (Neumann et al., 2024), climate change (Santurkar et al., 2023), and politics (Taubenfeld et al., 2024).
 In our study, we found that participants intentionally assessed whether LLM simulations accurately reflected the roles they represented (P19, 29, 31, 34, 38-39) and revised the descriptions if they considered simulations might be biased (P29, 34).
@@
 P2 took on the persona of an animal to test if the LLM could effectively respond from an animal perspective on animal testing “I am surprised by the cat persona, it actually understands the description, hates dogs, and says things less logically.
-Of course, users may hold their own biases toward certain social roles which could further reinforce their confirmation biases.
-The very flexibility our system enables users to customize the LLM personas and correct biases they find also means that they have the power to customize personas that reaffirm their existing beliefs.
+Of course, users may hold their own biases toward certain social roles, which could further reinforce their confirmation biases.
+The flexibility of our system enables users to customize the LLM personas and correct biases they find, which also means that they have the power to customize personas that reaffirm their existing beliefs.
 We cannot objectively evaluate how well participant customization reflects social roles within the demographic groups embodied by those personas.
@@
 Additionally, HCI research might consider how to balance giving users control over customization with ensuring representational diversity and fairness in LLM-generated personas.
-6.4. Reducing Automation Bias in AI-Assisted Decision-Making
-Although we did not explicitly assess the system’s impact on automation bias (Bertrand et al., 2022) (i.e., whether participants would automatically follow AI suggestions or engage in deliberate decision-making), the debate format of arguing opposing positions directly avoids providing simple answers to users that might be accepted without deliberate engagement.
-The debate format, with high customization, forces users to think critically and objectively assess the different arguments presented.
+6.4. Addressing Automation Bias in AI-Assisted Decision-Making
+While our study primarily examines the use of LLM multi-personas in reducing confirmation bias, other studies have explored similar applications in the context of AI-assisted decision-making.
+For example, Chiang et al.
+(2024) found that LLM could play as a devil advocate to foster user appropriate reliance on AI decisions.
+This suggests that LLM-driven persona debates may also help mitigate automation bias (Bertrand et al., 2022) (i.e., whether participants would automatically follow AI suggestions or engage in deliberate decision-making).
+In this section, we discuss relevant insights from participant feedback on this potential benefit.
+The debate format of arguing opposing positions directly avoids providing simple answers to users that might be accepted without deliberate engagement.
+Its high level of customization encourage users to think critically and objectively assess the different arguments presented.
 This contrasts with the more passive role users typically take in traditional AI-assisted decision-making scenarios, where they might simply accept AI decisions (Buçinca et al., 2021; Lai et al., 2023).
-Many participants intentionally structured a robust debate by selecting or customizing personas with different stances on the question (P14, 22, 24).
+For example, many participants intentionally structured a robust debate by selecting or customizing personas with different stances on the question (P14, 22, 24).
 P22 stated, “I felt like the descriptions that it gave would give a good debate” P24 added, “I’d probably change these personas a little more into what I feel like would really structure a good debate.” Although asking people to grapple with uncertainty may be uncomfortable, this encourages users to engage with multiple viewpoints and think critically to promote more deliberate decision-making.
@@
 6.5.1. System design and development
-As described in Section 6.1, our current multi-persona debate system lacks some interactive features to increase user engagement, such as the ability to nudge debate conversations by incorporating user feedback and new inputs.
+Our current multi-persona debate system lacks some interactive features to increase user engagement, such as the ability to nudge debate conversations by incorporating user feedback and new inputs.
 Additionally, system improvements are needed to enhance debate quality (Section 6.2), such as integrating RAG to reduce potential hallucination problems, refining prompting techniques to minimize repetitive responses from personas, and facilitating a natural closure in the debate (e.g., “I have nothing new to add”).
+While our system was built on top of ArgumentSearch to enable viewpoint diversity and did not incorporate traditional search engines as another baseline for comparison given the burden of evidence of confirmation bias observed (Section 2.2), it is important to acknowledge that this choice may limit the generalizability of our findings to more typical search scenarios.
+Future studies could include comparative evaluations with traditional or commercial search engines to further assess external validity and robustness.
 Additionally, because the system has not been evaluated for potential biases in its representations across topics and the persona, it is highly likely that biases in the LLM could impact the results of our user testing.
@@
 When a user enters a query, such as “Do violent video games contribute to youth violence?” the system identifies relevant articles and presents content snippets representing both sides, including those in favor (PRO) and those opposed (CON) to the topic.
+[…]
+Screenshot of the baseline retrieval system
 ArgumentSearch is a retrieval-based search engine that identifies articles from the open web and organizes content snippets according to their stance on the user query.
@@
 Table 5. Participant information
-Appendix E Additional Findings regarding User Behavior Patterns
-Beyond the main behavior clusters outlined in Section 5.1, in this section, we provide additional analysis, describing the topical differences within these groups (Section E.1) and additional qualitative findings regarding how our participants select and revise personas (Section E.2).
+Appendix E Additional Findings
+Beyond the main behavior clusters and qualitative themes outlined in Section 5.1, we present additional quantitative findings.
+These include participants’ behavioral patterns across different topics (Section E.1) their descriptive behavior data (Section E.2), detailed generalized linear mixed-effects model analysis (Section E.3), and user distribution of bi-directional belief change across topics (Section E.4).
 E.1. Topical differences among user behavioral clusters
 Figure 9. Distribution of user behavior patterns across different topics.
+[…]
+Distribution of user behavior patterns across different topics
 Additionally, when we separated these behavioral groups across different topics, we observed significant differences.
@@
 Overall, these results suggest that topic-specific factors do influence participant engagement patterns in using the debate system.
-E.2. Persona expertise, individuality, and diversity in selecting and revising personas
-From the post-task interviews, we found three additional factors that participants frequently mentioned in selecting and revising personas, including persona expertise, individuality, and the overall diversity of personas when selecting and editing them to construct the debate.
-These factors were closely tied to how participants conceptualized the controversial topics being discussed.
-First, many participants preferred selecting both expert and individual personas.
-Expert personas were more trusted to provide factual information, evidence, and other logical and reliable arguments to the controversial topic (P11, 14-15, 22, 24, 33, 36, 38-39).
-P39 explained that for the topic of “TikTok ban,” experts could provide opinions rather than emotional responses “Parents who are concerned about their kid is great, but I don’t expect them to have all of the data necessary to make an informed decision.
-I thought there would be experts discussing things logically and providing evidence-based information, as opposed to information driven by emotions.” In contrast, individual personas could provide direct and authentic arguments (P11, 22, 24-25, 33, 38).
-For example, P38 said “Whether TikTok should be banned from school.
-I really want from people who are in the school specifically because it’s important to see the people who are actually being affected to know their pain points.
-So that’s why I picked the teacher, the student, and the administrator.
-Additionally, as participants were aware that potential bias could underlie different personas, participants aimed to create a balanced representation of opinions by selecting personas that supported different sides of the topic.
-On one hand, this approach could be an effective strategy for generating counterarguments (P14, 24, 26, 28, 33, 35-36), as P36 explained, “I tried to balance the number of people who supported the topic with those who opposed it, to sort of weigh which perspective truly aligns with my own thinking and against it.” On the other hand, some would deliberately balance opinions that might be disproportionately represented on one side in real-world discussions (P22, 24, 33).
-P33 pointed out that“If I’ve been hearing a lot of one point of view, I’ll try to keep generating a persona until I find one that opposes it.
-Appendix F Additional Statistical Results
-F.1. Descriptive user behavior data in debate system
+E.2. Descriptive user behavior data in debate system
 Mean Std. Deviation Minimum Maximum
@@
 Table 6. Descriptive statistics of interactive behavior (mentioned in Section  4.4) in the multi-persona debate system.
-F.2. Generalized linear mixed-effects model analysis
+E.3. Generalized linear mixed-effects model analysis
 Variables Estimate Std. Error Z value P value
@@
 We code higher value for each variable: a higher score on CB and TF reflects increased tendency on confirmation bias and topic familiarity; the experiment condition is coded with a higher value than the baseline.
-F.3. Distribution of bi-directional belief changes across topics
+E.4. Distribution of directional belief changes across topics
 ID Topics Groups Bi-directional belief changes (# — %)
@@
 Table 9.
-Bi-directional belief changes across different topics: The retrieval-based and multi-persona systems have varying effects on user belief changes depending on the topic.
-For topics 1 and 2, the retrieval-based system tends to reinforce participants’ previous stances, making them more supportive (highlighted as orange).
-In contrast, for topics 5 and 8, the multi-persona debate system leads participants to become less supportive of their initial stance (highlighted as lightblue).
+Directional belief changes across different topics: The retrieval-based and multi-persona systems have varying effects on user belief changes depending on the topic.
+For topics 1 and 2, the retrieval-based system tends to reinforce participants’ supportive stance (highlighted as orange).
+In contrast, for topics 5 and 8, the multi-persona debate system leads participants to become less supportive (highlighted as lightblue).
 Both systems only show subtle impacts on other topics, but they both lead to participants becoming more supportive on topic 6 (highlighted as gray).
 By mapping participant bidirectional belief changes across different topics (Table 9), we found that the yielded system effects on weighted belief changes also vary on the topic.
-For example, in topics “Should animals be used for scientific or commercial testing?” and “Should humans colonize space?” (highlighted as orange) participants became more supportive of their previous stances after using the baseline.
-In contrast, for topics “Should the Federal corporate income tax rate be raised?” and “Do violent video games contribute to youth violence?” (highlighted as blue) participants became less supportive of their previous stance after using the experiment.
+For example, in topics ‘Animal Testing” and ‘Space Colonization’ (highlighted as orange) participants became more supportive of their previous stances after using the baseline.
+In contrast, for topics ‘Federal Corporate Income Tax Raise’ and ‘Violent Video Games Effect on Youth Violence’ (highlighted as blue) participants became less supportive of their previous stance after using the experiment.
 This suggests that user belief changes differ depending on the topic, with some topics being particularly sensitive.
+Reflecting on user qualitative insights as summarized in Section 5.1.2 and 5.1.3, we learned that topics like ‘Animal Testing’ and ‘Space Colonization’ may have been influenced by the retrieval system’s ability to present factual, research-driven information, reinforcing expert arguments in favor of these issues.
+In contrast, the interactive format of the debate system may have heightened participant exposure to counterarguments, leading to decreased support for policies that heavily influence each individual, such as ‘Federal Corporate Income Tax Raise’ and ‘Violent Video Games Effect on Youth Violence.’ For these topics, opposition arguments presented by the debate system tend to be more persuasive.
+Moreover, the overall increase in support for ‘Universal Basic Income’ across both systems suggests a general receptivity to economic policies that emphasize social welfare.